### v2

# Equivalence of Landscape and Erosion Distances for Persistence Diagrams

Abstract
This paper establishes connections between three of the most prominent metrics used in the analysis of persistence diagrams in topological data analysis: the bottleneck distance, Patel’s erosion distance, and Bubenik’s landscape distance. Our main result shows that the erosion and landscape distances are equal, thereby bridging the former’s natural category-theoretic interpretation with the latter’s computationally convenient structure. The proof utilizes the category with a flow framework of de Silva et al., and leads to additional insights into the structure of persistence landscapes. Our equivalence result is applied to prove several results on the geometry of the erosion distance. We show that the erosion distance is not a length metric, and that its intrinsic metric is the bottleneck distance. We also show that the erosion distance does not coarsely embed into any Hilbert space, even when restricted to persistence diagrams arising from degree-0 persistent homology. Moreover, we show that erosion distance agrees with bottleneck distance on this subspace, so that our non-embeddability theorem generalizes several results in the recent literature.
1 Introduction
Persistent homology is the most prominent tool in the field of Topological Data Analysis (TDA). In its classical form, persistent homology analyzes a one-parameter family of topological spaces (typically arising as a multiscale representation of a dataset) and outputs a topological summary known as a persistence diagram [48, 11]. Persistence diagrams take the form of multisets of points $(b,d)\in\mathbb{R}^{2}$ with $b<d$, each of which represents the lifespan of a homological feature in the input family of topological spaces. The space of persistence diagrams, which we denote $\mathsf{PDgm}$, can be endowed with a variety of metrics, including the bottleneck distance [20], Wasserstein distances [21], erosion distance [38], and various (pseudo)metrics arising from embedding diagrams into Hilbert or Banach spaces for machine learning tasks (see [15, 6, 1, 18], among many others). Within the TDA community, there has been significant recent interest in exploring properties of these metrics [8, 45, 37, 40, 16].
This paper is concerned with connections between three of the most widely used metrics on $\mathsf{PDgm}$. First, the bottleneck distance [20] is considered to be the canonical metric on $\mathsf{PDgm}$, due in part to its stability to noise in the input data and its natural interpretation in the language of category theory. Moreover, the bottleneck distance can be computed in polynomial time via combinatorial optimization (see Remark 4.14). Second, the landscape distance [6] arises via a certain embedding of $\mathsf{PDgm}$ into a Banach space, called the landscape map. This leads to a very easy-to-compute metric which is more amenable to statistical applications than the bottleneck distance, by performing them in the ambient Banach space. However, it does not have an immediate category-theoretical interpretation, and the image of the Banach space embedding is not well understood. Finally, the erosion distance [38] is an alternative metric on the space of persistence diagrams with category-theoretic origins. It provides a lower bound on bottleneck distance (hence enjoying the same stability to perturbations in the data), it has been suggested that it is efficient to compute (see the discussion surrounding [46, Example 1]), and it has found several uses in theory and applications, e.g., [30, 47, 31, 12].
Our main result shows that the latter two metrics discussed above are, in fact, the same:
Theorem (Theorem 3.19).
The landscape map is an isometry between the erosion distance and the landscape distance.
To prove the theorem, we adopt the category with a flow framework developed by de Silva et al. in [24], which involves realizing the erosion and landscape distances in this category-theoretic framework. While a more direct proof of this fact is possible—for instance, it can be derived from the more general result [31, Theorem 4.3] (see Remark 3.20)—we chose our particular approach because it sheds additional light on the structure of persistence landscapes. Indeed, in preparation for the proof of the main theorem, we provide in Theorem 3.13 a novel characterization of the image of the landscape map. This result, in turn, utilizes a new algebraic structure on the space of persistence landscapes (Theorem 3.17).
The main theorem directly leads to several new results on the metric geometry of the erosion distance and its connections to bottleneck distance. Our main contributions in this direction are as follows:
  • It is known that the erosion distance is not equal to the bottleneck distance, in general—in fact, we show in Proposition 4.1 that they are not even coarsely equivalent. However, we prove that they are locally equivalent; in particular, they generate the same topology on $\mathsf{PDgm}$ (Theorem 4.2). In fact, we show that the bottleneck distance is generated as the intrinsic (geodesic) metric associated to the erosion distance (Theorem 4.7). Accordingly, bottleneck distance can be seen as the intrinsic metric associated to landscape distance. This answers questions posed in [31, Example 3.15], on whether the erosion distance is geodesic.
  • We restrict our attention to a specific subspace of persistence diagrams, namely, birth-zero diagrams, denoted $\mathsf{PDgm}^{0}$ (persistence diagrams where each point is of the form $(0,d)$). We show that, on this subspace, the bottleneck and erosion (and hence the landscape) distances are equal (Theorem 4.12).
  • There is a simple map taking $\mathsf{PDgm}^{0}$ into a Banach space; namely, the death-vectorization embedding $\mathrm{DV}:\mathsf{PDgm}^{0}\to\ell^{\infty}$, introduced by Patrangenaru et. al. [39], sorts the non-zero coordinates of the points in a diagram to obtain a finitely-supported, non-increasing sequence. As a corollary to Theorem 4.12, we prove that this map is bi-Lipschitz with respect to the bottleneck/erosion/landscape distances (Corollary 4.18). This refines the main stability result of [36] (see Remark 4.20).
  • Finally, we leverage our results on $\mathsf{PDgm}^{0}$ to show that this space, endowed with bottleneck/erosion/landscape distance does not admit a coarse embedding into a Hilbert space. It follows that the full space $\mathsf{PDgm}$, endowed with any of the metrics we consider, also fails to admit such a coarse embedding (Theorem 4.21). Coarse embedding questions in TDA have seen quite a bit of activity in recent years, and our work generalizes the main results of [8, 37].
The structure of the paper is as follows. In Section 2, we provide the reader with the required background material on TDA invariants and metrics. Section 3 is devoted to proving our main theorem, Theorem 3.19, which says that erosion and landscape distances coincide. Our approach is to define categorical structures on the space of diagrams and on the space of landscapes, equip them with coflows, and show that the induced interleaving distances are equivalent, in the sense of [24]. The main theorem is applied in Section 4 to derive geometrical results on the erosion distance, where we shift focus on its connections to bottleneck distance and coarse embeddability questions.
2 Background
The focus of this paper is on connections between various invariants arising in the field of Topological Data Analysis (TDA). Throughout the paper, we assume that the reader is familiar with the basic concepts of TDA—see [10, 25] as general references. This section tersely recalls the main concepts of interest, and mostly serves to set terminology and notation. With this terminology in place, we also recapitulate some of the main results of the paper in precise language.
2.1 Topological Invariants of Data
We begin by reviewing some of the main dataset descriptors studied in TDA.
2.1.1 Persistence Modules
A persistence module is a functor
$\mathcal{F}:(\mathbb{R},\leq)\to\mathsf{vect}$
where $(\mathbb{R},\leq)$ is the poset category of $\mathbb{R}$, and $\mathsf{vect}$ is the category of finite-dimensional vector spaces over a fixed field $\mathbb{F}$. This means that:
  • To each $r\in\mathbb{R}$, we assign a vector space $\mathcal{F}(r)\in\mathsf{vect}$.
  • To each pair of real numbers $r\leq s$, we assign a linear map $\mathcal{F}(r\leq s):\mathcal{F}(r)\to\mathcal{F}(s)$.
satisfying the usual functorial properties: for any $r\in\mathbb{R}$, the map $\mathcal{F}(r\leq r)$ is the identity map $\mathcal{I}_{\mathcal{F}(r)}$ on $\mathcal{F}(r)$, and for any $r\leq s\leq t$, the composition rule $\mathcal{F}(r\leq t)=\mathcal{F}(s\leq t)\circ\mathcal{F}(r\leq s)$ holds.
In practice, persistence modules typically arise by applying the (degree-$k$) homology functor to a filtered simplicial complex; that is, to a functor from the poset category $(\mathbb{R},\leq)$ into $\mathsf{SpCpx}$, the category of finite simplicial complexes with simplicial maps. For example, such a structure arises as the Vietoris-Rips complex of a finite metric space. A persistence module of this form is referred to as the (degree-$k$) persistent homology of the filtered simplicial complex.
As they are functors, the appropriate notion of a morphism between persistence modules $\mathcal{F}$ and $\mathcal{G}$ is a natural transformation. This amounts to a family of linear maps $(\varphi_{r}:\mathcal{F}(r)\to\mathcal{G}(r))_{r\in\mathbb{R}}$ such that, whenever $r\leq s$,
$\mathcal{G}(r\leq s)\circ\varphi_{r}=\varphi_{s}\circ\mathcal{F}(r\leq s).$
Then an isomorphism is a morphism $(\varphi_{r})_{r}$ such that each $\varphi_{r}$ is an isomorphism of vector spaces.
2.1.2 Persistence Diagrams
A persistence module can be decomposed as a direct sum of 1-dimensional modules, each of which can be represented by a pair of (extended) real numbers $(b,d)$ with $\infty\leq b<d\leq\infty$ [22]. This decomposition gives rise to a persistence diagram, or a multiset $\{(b_{i},d_{i})\}$ of pairs with $b_{i}<d_{i}$.
Persistence diagrams are the main dataset invariant used in TDA. When a diagram arises as a representation of persistent homology, each point $(b,d)$ is interpreted as describing the “birth” and “death” scales of a homological feature in the filtered simplicial complex.
In this paper, we make simplifying assumptions that persistence diagrams are finite multisets, whose points satisfy $\infty<b<d<\infty$. This is partially for convenience and partially for technical reasons—see Remark 3.2 for details. In any case, this aligns with the structure of persistence diagrams in computational applications, where one necessarily works with finite collections and where points with $d=\infty$ are either ignored or truncated to some fixed, large but finite, value.
Let $\mathbb{R}^{2}_{<}=\{(b,d)\in\mathbb{R}\times\mathbb{R}\mid b<d\}$ denote the set of points in $\mathbb{R}^{2}$ above the diagonal. With the conventions described above, a persistence diagram in our setting is a multiset of the form $Y=\{(b_{i},d_{i})\}_{i=1}^{N}$, such that $(b_{i},d_{i})\in\mathbb{R}^{2}_{<}$ for each $i$.
2.1.3 Rank Functions
Let $Y=\{(b_{i},d_{i})\}_{i=1}^{N}$ be a persistence diagram. The associated rank function is the map
$\mathrm{rank}[Y]:\mathbb{R}^{2}_{\leq}\to\mathbb{Z},$
where $\mathbb{R}^{2}_{\leq}\coloneqq\{(b,d)\in\mathbb{R}^{2}\mid b\leq d\}$, defined by
$\mathrm{rank}[Y](b,d)\coloneqq|\{(b_{i},d_{i})\in Y\mid b_{i}\leq b\leq d<d_{i}\}|.$ (1)
In the setting of persistent homology, this function captures the number of topological features which persist over the interval $(b,d)$. Indeed, if $Y$ arises as a decomposition of a persistence module $\mathcal{F}$, then the rank can also be expressed as
$\mathrm{rank}[Y](b,d)=\mathrm{rank}(\mathcal{F}(b\leq d)),$ (2)
where the latter is the rank in the usual sense of linear algebra.
Terminology 2.1.
We say that $(b_{i},d_{i})\in Y$ properly contains $(b,d)\in\mathbb{R}^{2}_{\leq}$ if $b_{i}\leq b$ and $d<d_{i}$. With this terminology in place, $\mathrm{rank}[Y](b,d)$ is the number of pairs $(b_{i},d_{i})\in Y$ that properly contain $(b,d)$.
2.1.4 Persistence Landscapes
Persistence landscapes, introduced by Bubenik in [6], are a powerful representation of persistence diagrams. This stable representation has been shown to be both machine-learning friendly and statistically accessible [6, 4], as it provides a method to map persistence diagrams (which are somewhat unwieldy from a classical statistics perspective) to vectors in a Banach space.
Let $Y$ be a persistence diagram. As defined by Bubenik in [6], the persistence landscape of $Y$ is the function $\lambda^{Y}:\mathbb{N}\times\mathbb{R}\to\mathbb{R}$ defined by
$\lambda^{Y}(k,t)=\sup\{h\geq 0\mid\,\mathrm{rank}[Y](t-h,t+h)\geq k\}.$
Alternatively, $\lambda^{Y}$ can be viewed as a sequence of functions $\lambda^{Y}=(\lambda^{Y}_{1},\lambda^{Y}_{2},\dots)$, where $\lambda^{Y}_{k}\coloneqq\lambda^{Y}(k,\cdot)$ is called the $k$-th landscape function. As each landscape function is bounded, and only finitely many of them are not identically zero, $\lambda^{Y}$ lies in the Banach space $L^{\infty}(\mathbb{N}\times\mathbb{R})$.
In [4], Bubenik gives an equivalent definition of persistence landscapes in terms of the birth-death pairs, which we state below. First, for any $(b,d)\in\mathbb{R}^{2}_{<}$, define the associated tent function $f_{(b,d)}:\mathbb{R}\to\mathbb{R}$ by
$f_{(b,d)}(t)=\begin{cases}0,\text{ if }t\notin(b,d)\\ t-b,\text{ if }t\in(b,\frac{b+d}{2}]\\ d-t,\text{ if }t\in(\frac{b+d}{2},d).\end{cases}$ (3)
We then state the equivalent definition from [4] as a proposition.
Proposition 2.2 ([4]).
Let $Y=\{(b_{i},d_{i})\}_{i=1}^{N}$ be a persistence diagram. The associated $k$-th landscape function can be expressed as
$\lambda^{Y}_{k}(t)=k\mathrm{max}\{f_{(b_{i},d_{i})}(t)\,\mid\,i=1,2,\dots,N\}$
where $k\mathrm{max}$ stands for the $k$-th maximum element, which we declare to be 0 for $k>N$.
One of the main contributions of this paper is an exact characterization of the image of the map $Y\mapsto\lambda^{Y}$ taking a diagram to its landscape—see Theorem 3.13 and Corollary 3.14.
In Figure 1, we present an example of a persistence diagram $Y$, along with values of its rank function $\mathrm{rank}[Y]$, and its persistence landscape $\lambda^{Y}$. Note that, according to Theorem 3.17, $\lambda^{Y}$ can be expressed as
$\lambda^{Y}=f_{(1,7)}\oplus f_{(3,8)}\oplus f_{(2,5)}\oplus f_{(2,5)}\oplus f_{(9,10)}$
where $\oplus$ denotes the direct sum operation defined in Definition 3.15.
Figure 1: Left: An example persistence diagram $Y=\{(1,7),(3,8),(2,5),(2,5),(9,10)\}$. The light-gray lines extending horizontally and vertically from each birth-death pair reveal the formation of the persistence landscape. Moreover, the values of $\mathrm{rank}[Y]$ are annotated in blue, associated with each region formed by the underlying landscape. Right: The corresponding persistence landscape $\lambda^{Y}$ which forms a non-increasing sequence of piecewise-linear functions $\lambda^{Y}_{1}\geq\lambda^{Y}_{2}\geq\lambda^{Y}_{3}\geq\lambda^{Y}_{4}$ (and $\lambda^{Y}_{k}=0$ for $k\geq 5$), as indicated by different colors.
2.2 Standard Metrics in TDA
We now overview some metrics used to compare the TDA invariants described in the previous subsection.
2.2.1 Bottleneck Distance
Let $Y=\{(b_{i},d_{i})\}_{i=1}^{N}$ and $Y^{\prime}=\{(b_{i}^{\prime},d_{i}^{\prime})\}_{i=1}^{N^{\prime}}$ be two persistence diagrams. The bottleneck distance [20] between $Y$ and $Y^{\prime}$ is given by
$d_{B}(Y,Y^{\prime})\coloneqq\min_{\phi:S\to S^{\prime}}\max\left\{\max_{(b,d)\in S}d_{\infty}((b,d),\phi(b,d)),\;\max_{(b,d)\notin S}\frac{(d-b)}{2},\;\max_{(b^{\prime},d^{\prime})\notin S^{\prime}}\frac{(d^{\prime}-b^{\prime})}{2}\right\}$
where
  • $\phi:S\to S^{\prime}$ is a bijection with $S\subseteq Y$ and $S^{\prime}\subseteq Y^{\prime}$,
  • $d_{\infty}((b,d),(b^{\prime},d^{\prime}))\coloneqq\max\{|b-b^{\prime}|,|d-d^{\prime}|\}$.
Intuitively, $d_{B}(Y,Y^{\prime})$ finds a matching between the topological features corresponding to points in the diagrams, which is optimal with respect to the $\ell^{\infty}$-type cost. The bottleneck distance is known to be stable to perturbations in the data from which the diagrams are derived [19], and its computation is a tractable optimization problem; these properties, in part, lead to its status as the de facto canonical metric on the space of persistence diagrams.
2.2.2 Interleaving Distance
Loosely speaking, one can define a metric on the space of persistence modules by quantifying the degree to which any pair is non-isomorphic; this motivates the concept of interleaving distance, as introduced by Chazal et al. [13], which we now recall. For $\varepsilon\geq 0$, two persistence modules $\mathcal{F}$ and $\mathcal{G}$ are said to be $\varepsilon$-interleaved if, for all $r\in\mathbb{R}$, there are linear maps $\varphi_{r}:\mathcal{F}(r)\to\mathcal{G}(r+\varepsilon)$ and $\psi_{r}:\mathcal{G}(r)\to\mathcal{F}(r+\varepsilon)$ such that the following diagrams commute for any $r\leq s$:
$\mathcal{F}(r)$$\mathcal{F}(s)$$\mathcal{F}(r+\varepsilon)$$\mathcal{F}(s+\varepsilon)$$\mathcal{G}(r+\varepsilon)$$\mathcal{G}(s+\varepsilon)$$\mathcal{G}(r)$$\mathcal{G}(s)$$\mathcal{F}(r)$$\mathcal{F}(r+2\varepsilon)$$\mathcal{F}(r+\varepsilon)$$\mathcal{G}(r+\varepsilon)$$\mathcal{G}(r)$$\mathcal{G}(r+2\varepsilon)$$\scriptstyle{\mathcal{F}(r\leq s)}$$\scriptstyle{\varphi_{r}}$$\scriptstyle{\varphi_{s}}$$\scriptstyle{\mathcal{F}(r+\varepsilon\leq s+\varepsilon)}$$\scriptstyle{\mathcal{G}(r+\varepsilon\leq s+\varepsilon)}$$\scriptstyle{\psi_{r}}$$\scriptstyle{\mathcal{G}(r\leq s)}$$\scriptstyle{\psi_{s}}$$\scriptstyle{\mathcal{F}(r\leq r+2\varepsilon)}$$\scriptstyle{\varphi_{r}}$$\scriptstyle{\varphi_{r+\varepsilon}}$$\scriptstyle{\psi_{r+\varepsilon}}$$\scriptstyle{\psi_{r}}$$\scriptstyle{\mathcal{G}(r\leq r+2\varepsilon)}$
The interleaving distance between $\mathcal{F}$ and $\mathcal{G}$ is then defined to be
$d_{I}(\mathcal{F},\mathcal{G})=\inf\{\varepsilon\geq 0\mid\mbox{$\mathcal{F}$ and $\mathcal{G}$ are $\varepsilon$-interleaved}\}.$
In general, the interleaving distance defines an extended pseudometric on the space of persistence modules. In fact, a fundamental theorem in TDA, referred to as the isometry theorem [13, 33], says that the interleaving distance between persistence modules is equal to the bottleneck distance between their persistence diagram representations.
The notion of interleaving distance has been extended over the years to define metrics for increasingly general structures [7, 5, 24, 44, 43, 34]. An appropriate level of generality for our purposes is provided by the category with a flow framework of de Silva et al. [24], as we explain below in Section 2.3.
2.2.3 Erosion Distance
Due to its stability properties and its connection to the interleaving distance (as described above), the bottleneck distance $d_{B}$ is considered to be the canonical metric on the space of persistence diagrams. However, the erosion distance, introduced by Patel in [38], is also motivated by category theory and enjoys stability to perturbations in the data. We present a definition from [46] here, and give Patel’s (equivalent) definition below in Section 3.1.2. The erosion distance between two persistence diagrams $Y$ and $Y^{\prime}$ is given by
$\begin{split}d_{E}(Y,Y^{\prime})\coloneqq\inf\{\varepsilon\geq 0\mid\mathrm{rank}&[Y](b-\varepsilon,d+\varepsilon)\leq\mathrm{rank}[Y^{\prime}](b,d)\\ &\mbox{ and }\mathrm{rank}[Y^{\prime}](b-\varepsilon,d+\varepsilon)\leq\mathrm{rank}[Y](b,d)\mbox{ for all }b\leq d\}.\end{split}$ (4)
It is not hard to show that $d_{E}\leq d_{B}$, nor is it hard to construct an example which illustrates that this is not an equality, in general (see, e.g., [46, Example 1]). However, we prove in Theorem 4.2 that these metrics are locally equivalent, and therefore induce the same topology on the space of diagrams. Moreover, we show in Theorem 4.7 that the bottleneck distance is the intrinsic distance induced by the erosion distance.
2.2.4 Landscape Distance
As was observed in Section 2.1.4, the persistence landscape $\lambda^{Y}$ of a diagram $Y$ lies in the Banach space $L^{\infty}(\mathbb{N}\times\mathbb{R})$. Landscapes are then naturally compared by the metric induced by the norm, referred to as the landscape distance:
$\|\lambda^{Y}-\lambda^{Y^{\prime}}\|_{\infty}\coloneqq\sup_{k,t}|\lambda^{Y}_{k}(t)-\lambda^{Y^{\prime}}_{k}(t)|.$ (5)
The main result of this paper, Theorem 3.19, shows that the landscape map $Y\mapsto\lambda^{Y}$ is an isometry with respect to the erosion and landscape distances. This result is leveraged to prove that the space of diagrams, endowed with erosion distance, does not coarsely embed into a Hilbert space (Theorem 4.21).
2.3 Coflows on Poset Categories
We will make heavy use of the category with a flow framework of [24], which extends the interleaving distance construction (described in Section 2.2.2) to define a very general class of metrics, presented in the language of category theory. For our purposes, it suffices to specialize the framework to the setting of poset categories. Moreover, [24] distinguishes between lax and strict versions of the framework, whereas we only require the strict version. Finally, it turns out to be more convenient to use a related construction called a coflow, as introduced in [23]. For the sake of simplicity, we opt to present the framework specifically in the context of strict coflows on poset categories. For the rest of the paper, we frequently write $\mathsf{P}=(\mathsf{P},\leq)$ for a poset category—that is, we abuse notation and use $\mathsf{P}$ to denote both the category and the underlying set. We note that the following definitions could alternatively be expressed in the language of superlinear families of translations, coined in [5].
Definition 2.3.
Let $\mathsf{P}=(\mathsf{P},\leq)$ be a poset category. A family $\gamma=(\gamma_{\varepsilon})_{\varepsilon\geq 0}$ of endofunctors $\gamma_{\varepsilon}:\mathsf{P}\to\mathsf{P}$ is called a coflow on $\mathsf{P}$ if
  1. 1.
    $\gamma_{0}=\mathcal{I}_{\mathsf{P}}$, where $\mathcal{I}_{\mathsf{P}}$ is the identity endofunctor on $\mathsf{P}$;
  2. 2.
    for all $\varepsilon\geq 0$ and $x\in\mathsf{P}$, $\gamma_{\varepsilon}(x)\leq x$;
  3. 3.
    $\gamma_{\varepsilon_{1}+\varepsilon_{2}}=\gamma_{\varepsilon_{1}}\circ\gamma_{\varepsilon_{2}}$ for all $\varepsilon_{1},\varepsilon_{2}\geq 0$.
We will occasionally denote (by further abuse of notation) a poset category with a coflow as a triplet $\mathsf{P}=(\mathsf{P},\leq,\gamma)$. In the following, for $x\in\mathsf{P}$ and $\varepsilon,\varepsilon_{1},\varepsilon_{2}\geq 0$, we write $\gamma_{\varepsilon}x$ in place of $\gamma_{\varepsilon}(x)$, and $\gamma_{\varepsilon_{1}}\gamma_{\varepsilon_{2}}$ in place of $\gamma_{\varepsilon_{1}}\circ\gamma_{\varepsilon_{2}}$.
For $\varepsilon\geq 0$, we say that $x,y\in\mathsf{P}$ are $\varepsilon$-interleaved (with respect to $\gamma$) if $\gamma_{\varepsilon}x\leq y$ and $\gamma_{\varepsilon}y\leq x$. The interleaving distance (associated to $\gamma$) between $x$ and $y$, denoted $d_{\mathsf{P}}^{\gamma}(x,y)$, is the infimum of all $\varepsilon\geq 0$ such that $x$ and $y$ are $\varepsilon$-interleaved. That is,
$d_{\mathsf{P}}^{\gamma}(x,y)\coloneqq\inf\{\varepsilon\geq 0\,\mid\,\gamma_{\varepsilon}x\leq y\text{ and }\gamma_{\varepsilon}y\leq x\}.$
Remark 2.4.
A coflow on $(\mathsf{P},\leq)$ is a flow (in the language of [24]) on the opposite poset category $(\mathsf{P},\leq_{\mathrm{op}})$, where $x\leq_{\mathrm{op}}y$ if and only if $x\geq y$.
In light of Remark 2.4, the general result [24, Theorem 2.7] on interleaving distances associated to flows immediately implies the following.
Proposition 2.5.
The interleaving distance $d_{\mathsf{P}}^{\gamma}$ associated to a coflow $\gamma$ defines an extended pseudometric on $\mathsf{P}$.
Posets with coflows are related by the following construction.
Definition 2.6.
Let $(\mathsf{C},\leq_{\mathsf{C}},\tau)$ and $(\mathsf{D},\leq_{\mathsf{D}},\gamma)$ be two poset categories with coflows. A functor $\mathcal{H}:\mathsf{C}\to\mathsf{D}$ is called coflow equivariant if $\mathcal{H}\tau_{\varepsilon}=\gamma_{\varepsilon}\mathcal{H}$. If, in addition, $\mathcal{H}$ is an equivalence of categories, then $\mathcal{H}$ is called a coflow equivariant equivalence and we say that $\mathsf{C}$ and $\mathsf{D}$ are coflow equivalent.
A coflow equivariant equivalence defines an isometry of the associated interleaving distances. This is stated precisely in the following result, whose proof is straightforward; it also follows directly from the more general result [24, Theorem 4.3].
Proposition 2.7.
Let $(\mathsf{C},\leq_{\mathsf{C}},\tau)$ and $(\mathsf{D},\leq_{\mathsf{D}},\gamma)$ be two poset categories with coflows and let $\mathcal{H}:\mathsf{C}\to\mathsf{D}$ be a coflow equivariant equivalence. Then, for any $a,b\in\mathsf{C}$, we have $d_{\mathsf{C}}^{\tau}(a,b)=d_{\mathsf{D}}^{\gamma}(\mathcal{H}(a),\mathcal{H}(b))$.
3 Category Structure of Persistence Diagrams and Landscapes
This section is dedicated to developing formal category structures on spaces of objects which are fundamental to the theory of topological data analysis (TDA), introduced in Section 2.
3.1 Persistence Diagrams
Combining results and constructions which already exist in the literature, we now describe the space of persistence diagrams (see Section 2.1.2) as a poset category, endowed with a coflow and associated interleaving distance, and show that this distance recovers Patel’s erosion distance [38] (see Section 2.2.3).
3.1.1 Poset Category of Persistence Diagrams
Recall that $\mathbb{R}^{2}_{<}=\{(b,d)\in\mathbb{R}\times\mathbb{R}\mid b<d\}$ denotes the set of points in $\mathbb{R}^{2}$ above the diagonal, and that $\mathbb{R}^{2}_{\leq}$ is the superset which also includes the diagonal. We frequently denote elements of $\mathbb{R}^{2}_{<}$ as $I=(b,d)$ (intended to evoke the idea that $I$ is an interval), which we refer to as birth-death pairs, and refer to the first coordinate as the birth and the second coordinate as the death of $I$. We equip $\mathbb{R}^{2}_{<}$ with the poset relation given by inclusion:
$(b,d)\subseteq(b^{\prime},d^{\prime})\iff b^{\prime}\leq b\text{ and }d\leq d^{\prime}.$
Finally, recall that a persistence diagram is a finite multiset of points in $\mathbb{R}^{2}_{<}$, which we typically denote as $Y=\{(b_{i},d_{i})\}_{i=1}^{N}$ (mildly abusing notation and eschewing multiset notation). As was stated in the introduction, we denote the collection of all persistence diagrams as $\mathsf{PDgm}$.
Remark 3.1 (Generalized Persistence Diagrams).
Patel  [38] considered generalized persistence diagrams as maps $Y:\mathbb{R}^{2}_{<}\to\mathsf{G}$ with finite support, where $\mathsf{G}$ is an abelian group. In the case that $\mathsf{G}=\mathbb{Z}$, and with the additional requirement that the map takes non-negative values, this is equivalent to our notion of a persistence diagram.
Remark 3.2 (Finiteness Assumptions).
One could study persistence diagrams with countably many points. For example, Bubenik and Wagner in  [8] defined persistence diagrams as functions $D:\mathcal{I}\to\mathbb{R}^{2}_{<}$, where $\mathcal{I}$ is a countable indexing set. Throughout this paper, we define persistence diagrams to be finite, following  [38], because the definition of Patel’s erosion distance relies on rank functions, which are integer-valued.
Frequently in TDA, one considers persistence diagrams where death coordinates are allowed to take the value $\infty$. Points with this property represent topological features which never vanish in the filtration. We explicitly work with finite birth-death values rather than allowing birth-death pairs with infinite persistence. This assumption is only for convenience, as it avoids the need for cumbersome case distinctions in various definitions and results. Nevertheless, all results remain valid if one chooses to include birth-death pairs with infinite persistence, with only some adjustments to the definitions that follow.
A poset structure on $\mathsf{PDgm}$ is defined in terms of the rank function (see Section 2.1.3) as follows.
Definition 3.3 ([38]).
We define the persistence diagram poset structure $\leq_{\mathsf{PDgm}}$ on $\mathsf{PDgm}$ as follows:
$Y\leq_{\mathsf{PDgm}}Y^{\prime}\iff\mathrm{rank}[Y](I)\leq\mathrm{rank}[Y^{\prime}](I)\text{ for all }I\in\mathbb{R}^{2}_{\leq}.$
As the context is always clear, we simplify the notation and write $Y\leq Y^{\prime}$ instead of $Y\leq_{\mathsf{PDgm}}Y^{\prime}$.
3.1.2 Erosion Distance is a Coflow-Induced Interleaving Distance
We now reformulate the erosion distance (4) in more categorical terms, in order to take advantage of the coflow framework of Section 2.3. This follows the original construction of Patel, translated into our notation.
Definition 3.4 ([38]).
Given $\varepsilon\geq 0$, we define the $\varepsilon$-grow map $\mathrm{Grow}_{\varepsilon}:\mathbb{R}_{\leq}^{2}\to\mathbb{R}_{\leq}^{2}$ by
$\mathrm{Grow}_{\varepsilon}(b,d)\coloneqq(b-\varepsilon,d+\varepsilon).$
The $\varepsilon$-grow map induces a functor $\nabla_{\varepsilon}:\mathsf{PDgm}\to\mathsf{PDgm}$ defined by
$\nabla_{\varepsilon}Y\coloneqq\{(b,d)\in\mathbb{R}^{2}_{<}\mid\mathrm{Grow}_{\varepsilon}(b,d)\in Y\}.$
The erosion distance, $d_{E}$, between $Y$ and $Y^{\prime}$ is defined by
$d_{E}(Y,Y^{\prime})\coloneqq\inf\{\varepsilon\geq 0\,|\,\nabla_{\varepsilon}Y^{\prime}\leq Y\text{ and }\nabla_{\varepsilon}Y\leq Y^{\prime}\}.$
We will now show that the family $\nabla=\{\nabla_{\varepsilon}\}_{\varepsilon\geq 0}$ is a coflow and that the corresponding interleaving distance is the erosion distance. This perspective on erosion distance was already explored by Puuska in [41], albeit in the language of superlinear families of translations [5], rather than that of flows on categories [24], and by Kim, Mémoli and Stefanou in [31]. We provide the details in our case of interest, for the convenience of the reader. The first step is to prove the following.
Proposition 3.5.
The family $\nabla=(\nabla_{\varepsilon})_{\varepsilon\geq 0}$ defines a coflow on $\mathsf{PDgm}$.
The proof will use the following results; the first two are obvious, while the last follows from [38, Proposition 5.1], but is proved here, in order to mitigate difficulties in translating between notational conventions.
Lemma 3.6.
For $\varepsilon_{1},\varepsilon_{2}\geq 0$, we have that $\mathrm{Grow}_{\varepsilon_{1}}\circ\mathrm{Grow}_{\varepsilon_{2}}=\mathrm{Grow}_{\varepsilon_{1}+\varepsilon_{2}}$.
Lemma 3.7.
The rank function $\mathrm{rank}[Y]$ is order-reversing. That is, if $I_{1}\subseteq I_{2}$, then
$\mathrm{rank}[Y](I_{2})\leq\mathrm{rank}[Y](I_{1}).$
Lemma 3.8 ([38, Proposition 5.1.]).
Let $Y$ be a persistence diagram. Then, given any $\varepsilon\geq 0$, we have $\mathrm{rank}[Y]\left(\mathrm{Grow}_{\varepsilon}(I)\right)=\mathrm{rank}[\nabla_{\varepsilon}Y](I)$ for all $I\in\mathbb{R}_{\leq}^{2}$.
Proof.
For $I=(b,d)\in\mathbb{R}_{\leq}^{2}$, we have
$\mathrm{rank}[Y]\left(\mathrm{Grow}_{\varepsilon}(I)\right)$ $=|\{(b_{i},d_{i})\in Y\mid b_{i}\leq b-\varepsilon\leq d+\varepsilon<d_{i}\}|$
$=|\{(b_{i},d_{i})\in Y\mid b_{i}+\varepsilon\leq b\leq d<d_{i}-\varepsilon\}|$
$=|\{(b_{i}+\varepsilon,d_{i}-\varepsilon)\in\mathbb{R}_{\leq}^{2}\mid(b_{i},d_{i})\in Y,\;b_{i}+\varepsilon\leq b\leq d<d_{i}-\varepsilon\}|$
$=|\{(b_{i}^{\prime},d_{i}^{\prime})\in\mathbb{R}_{\leq}^{2}\mid\mathrm{Grow}_{\varepsilon}(b_{i}^{\prime},d_{i}^{\prime})\in Y,\;b_{i}^{\prime}\leq b\leq d<d_{i}^{\prime}\}|$
$=\mathrm{rank}[\nabla_{\varepsilon}Y](I).$
Proof of Proposition 3.5.
Let $Y\in\mathsf{PDgm}$. First, observe that
$\nabla_{0}Y=\{(b,d)\in\mathbb{R}^{2}_{<}\mid\mathrm{Grow}_{0}(b,d)\in Y\}=Y,$
so that the first axiom in Definition 2.3 is satisfied. To prove that the second axiom holds, observe that, by Lemmas 3.7 and 3.8, we have, for all $I\in\mathbb{R}_{\leq}^{2}$,
$\mathrm{rank}[\nabla_{\varepsilon}Y](I)=\mathrm{rank}[Y](\mathrm{Grow}_{\varepsilon}(I))\leq\mathrm{rank}[Y](I).$
Thus $\nabla_{\varepsilon}Y\leq Y$. Finally, let $\varepsilon_{1},\varepsilon_{2}\geq 0$ and apply Lemma 3.6 to deduce that
$\begin{split}\nabla_{\varepsilon_{1}}\nabla_{\varepsilon_{2}}Y&=\{(b,d)\in\mathbb{R}^{2}_{<}\mid\mathrm{Grow}_{\varepsilon_{1}}\circ\mathrm{Grow}_{\varepsilon_{2}}(b,d)\in Y\}\\ &=\{(b,d)\in\mathbb{R}^{2}_{<}\mid\mathrm{Grow}_{\varepsilon_{1}+\varepsilon_{2}}(b,d)\in Y\}=\nabla_{\varepsilon_{1}+\varepsilon_{2}}Y.\end{split}$
This proves that the final axiom of a coflow is satisfied. ∎
This leads immediately to the main result of this subsection.
Proposition 3.9.
The interleaving distance induced by the coflow $\nabla$ is the erosion distance.
Proof.
This is clear by definition, as
$d_{\mathsf{PDgm}}^{\nabla}(Y,Y^{\prime})=\inf\{\varepsilon\geq 0\,|\,\nabla_{\varepsilon}Y\leq Y^{\prime}\text{ and }\nabla_{\varepsilon}Y^{\prime}\leq Y\}=d_{E}(Y,Y^{\prime}).$
3.2 Persistence Landscapes
We introduced the persistence landscape $\lambda^{Y}$ of a persistence diagram in Section 2.1.4. In this section, we derive a characterization of the image of the map $Y\mapsto\lambda^{Y}$ as a subset of $L^{\infty}(\mathbb{N}\times\mathbb{R})$. This is used in Section 3.3.1 to define a categorical structure on this subspace and to then interpret the landscape distance (5) in terms of a coflow.
3.2.1 Persistence Landscapes and Landscape Sequences
Our first goal is to define landscapes as abstract objects, apart from arising solely as invariants of persistence diagrams. We begin with a preliminary definition.
Definition 3.10.
A function $\mathbb{R}\to\mathbb{R}$ is called a landscape curve if:
  1. 1.
    It is non-negative, continuous and piecewise-linear.
  2. 2.
    It is compactly supported.
  3. 3.
    It is differentiable except at finitely many points, called critical points.
  4. 4.
    At non-critical support points, its derivative is +1+1 or 1-1.
These conditions imply that any critical point must be a local maximum or local minimum (exclusively) of the function.
We now define our abstract notion of a persistence landscape, referred to as a landscape sequence.
Definition 3.11.
A landscape sequence is a sequence $\lambda=(\lambda_{1},\lambda_{2},\dots)$ of functions $\lambda_{k}:\mathbb{R}\to\mathbb{R}$ satisfying:
  1. 1.
    (Structure of $\lambda_{k}$) Each $\lambda_{k}$ is a landscape curve.
  2. 2.
    (Monotonicity) For all $k\in\mathbb{N}$ and $t\in\mathbb{R}$, $\lambda_{k}(t)\geq\lambda_{k+1}(t)$.
  3. 3.
    (Finite support) There exists $K\in\mathbb{N}$ such that $\lambda_{k}=0$ for all $k\geq K$.
  4. 4.
    (Non-negative degree) Any point $(t,h)\in\mathbb{R}^{2}$ with $h>0$ has non-negative degree, where the degree of a point $(t,h)$ is defined by
[…]
Terminology 3.12.
The term critical point may refer either to a pair $(t,h)$ or to the coordinate $t$ alone; i.e., when the height $h$ is implicit or can be inferred from the function values, we may simply refer to $t$ as a critical point. The set of critical points of the landscape sequence $\lambda$ is defined to be the union of the sets of critical points of each $\lambda_{k}$.
With this abstract definition of landscape sequences in hand, the main result of this subsection is that landscape sequences are, in fact, in one-to-one correspondence with persistence landscapes (arising from a persistence diagram).
Theorem 3.13 (Structure of Persistence Landscapes).
The map $\mathsf{PDgm}\to L^{\infty}(\mathbb{N}\times\mathbb{R})$ given by $Y\mapsto\lambda^{Y}$ defines a bijection between the set of persistence diagrams and the set of landscape sequences. That is:
  1. 1.
    Given $Y\in\mathsf{PDgm}$, its persistence landscape $\lambda^{Y}$ is a landscape sequence.
  2. 2.
    Conversely, every landscape sequence $\lambda$ is a persistence landscape of a unique persistence diagram $Y\in\mathsf{PDgm}$.
The “forward” direction, part 1, is relatively straightforward, via structural properties of landscape functions that were established in [2, Section 5]. The converse, part 2, is less clear, and is the main focus of the proof. An immediate corollary is that we can explicitly describe the image of the persistence landscape map. As far as we are aware, such a description has not been previously established in the literature.
Corollary 3.14.
The image of the landscape map $Y\mapsto\lambda^{Y}$ is exactly the set of landscape sequences, which is a proper, nonlinear (moreover, non-affine linear), subset of the space of functions $L^{\infty}(\mathbb{N}\times\mathbb{R})$.
The proof of Theorem 3.13 will be derived from a preliminary result, concerning an algebraic description of the structure of landscape sequences. This description and the ensuing proof of Theorem 3.13 are provided in the following subsection.
3.2.2 Decomposition Theorem for Landscape Sequences
To describe the structure of the space of landscape sequences, we introduce the following algebraic operation. It uses the kmax function, introduced in Proposition 2.2.
Definition 3.15.
Given two landscape sequences $\eta$ and $\mu$, we define their direct sum
$\lambda=\eta\oplus\mu$
to be a sequence of functions $\lambda_{k}$ given by
$\lambda_{k}(t)=k\mathrm{max}\{\eta_{1}(t),\dots,\eta_{k}(t),\mu_{1}(t),\dots,\mu_{k}(t)\}.$
Proposition 3.16.
The sequence of functions defined by $\lambda=\eta\oplus\mu$ is also a landscape sequence.
Proof.
Let $\lambda=\eta\oplus\mu$, where $\eta$ and $\mu$ are two landscape sequences. We will omit the verification of Properties 1, 2, and 3 in Definition 3.11, as they follow directly from the construction of the direct sum operation and are relatively straightforward to verify. Our goal is to show that $\lambda$ satisfies Property 4.
Let $\{(t_{\ell},h_{\ell})\}$ denote the multiset of local minima of $\lambda$ with positive second coordinates, $h_{\ell}>0$. Such a critical point $(t_{\ell},h_{\ell})$ of $\lambda_{k}$ can arise in exactly one of four ways:
  1. 1.
    as a local minimum of some $\eta_{i}$;
  2. 2.
    as a local minimum of some $\mu_{i}$;
  3. 3.
    there are $\eta_{i}$, with negative slope at $t_{\ell}$, and $\mu_{j}$, with positive slope at $t_{\ell}$, such that, for $t$ in a sufficiently small neighborhood of $t_{\ell}$, $\lambda_{k}(t)=\eta_{i}(t)$ for $t<t_{\ell}$ and $\lambda_{k}(t)=\mu_{j}(t)$ for $t>t_{\ell}$; or,
  4. 4.
    the same as the above, with the roles of $\eta_{i}$ and $\mu_{j}$ switched: there are $\eta_{i}$ (positive slope at $t_{\ell}$) and $\mu_{j}$ (negative slope at $t_{\ell}$) such that, for $t$ in a sufficiently small neighborhood of $t_{\ell}$, $\lambda_{k}(t)=\mu_{i}(t)$ for $t<t_{\ell}$ and $\lambda_{k}(t)=\eta_{j}(t)$ for $t>t_{\ell}$.
The collection $\{(t_{\ell},h_{\ell})\}$ can be decomposed into four disjoint (possibly empty) multisets, according to the four possibilities described above. We will show that for each of these multisets, there is an injective map into the multiset of local maxima for $\lambda$. This means that, at each critical point, the number of times that it is a local maximum must be greater than or equal to the number of times that it is a local minimum; that is, its degree is non-negative.
Consider the sub-multiset of local minima arising as in case 1 (local minima of $\eta$). Since $\eta$ is itself a landscape sequence, there must be an injective map into the multiset of local maxima of $\eta$. We now observe that each such local maximum of $\eta$ (i.e., those that are at a local minimum of $\lambda$) injectively maps into a local maximum of $\lambda$. Indeed, let $m$ be the maximum index such that $\lambda_{m}(t_{\ell})=h_{\ell}$. It is not hard to show that $\lambda_{m}$ forms a local maximum at $(t_{\ell},h_{\ell})$. Moreover, if $\eta$ has another such local maximum, then we can let $m_{2}$ be the second-largest index such that $\lambda_{m_{2}}(t_{\ell})=h_{\ell}$; this too forms a local maximum at $(t_{\ell},h_{\ell})$, verifying injectivity. So we have constructed the desired injective map. The map is constructed similarly for case 2.
For each local minimum $(t_{\ell},h_{\ell})$ arising as in case 3, there is an associated local maximum described as follows (using the same indices from the description of case 3 above). For some $m>k$, there is a local maximum of $\lambda_{m}$ at $(t_{\ell},h_{\ell})$ where, for $t$ in a sufficiently small neighborhood of $t_{\ell}$, $\lambda_{m}(t)=\mu_{j}(t)$ for $t<t_{\ell}$ and $\lambda_{m}(t)=\eta_{i}(t)$ for $t>t_{\ell}$. This correspondence between local minima and associated local maxima gives the desired injective map, and case 4 is similar. ∎
Proposition 3.16 says that direct sum gives a well-defined binary operation on the space of landscape sequences. Next, we claim that any landscape sequence $\lambda$ can be expressed as a direct sum of tent functions, which are the simplest landscape sequences. To clarify this statement, let $b<d$ with associated tent function $f_{(b,d)}$ (see (3)). By abuse of notation we denote the associated landscape sequence by $f_{(b,d)}$, where this really means the sequence of functions
$(f_{(b,d)},0,0,\ldots).$
Theorem 3.17 (Landscape Decomposition Theorem).
Let $\lambda$ be a landscape sequence. There exists a finite multiset of tent functions $\{f_{(b_{i},d_{i})}\}$ such that
$\lambda=\bigoplus_{i}f_{(b_{i},d_{i})},$
where each summand $f_{(b_{i},d_{i})}$ appears with multiplicity
$\mathrm{deg}_{\lambda}\bigg{(}\frac{b_{i}+d_{i}}{2},\frac{d_{i}-b_{i}}{2}\bigg{)}>0.$
Moreover, this decomposition is unique up to ordering of summands.
Before we prove Theorem 3.17, we provide a Lemma observing the effect of removing a particular tent function from a landscape sequence. In the lemma, and through the rest of the paper, for a landscape curve $\lambda_{i}$, we use $D^{-}\lambda_{i}(t)$ (respectively, $D^{+}\lambda_{i}(t)$) to denote the derivative of $\lambda_{i}$ from the left (respectively right) at the point $t$; this is well-defined at every point for a landscape curve.
Lemma 3.18.
Let $\lambda$ be a landscape sequence, and assume that $(x_{1},y_{1})$ is the left-most point where $\lambda_{1}$ attains a local maximum. Consider points
$x_{1}\leq x_{2}\leq x_{3}\leq\dots\leq x_{n}$
where each $x_{i}$, with $1\leq i\leq n$, is paired with $y_{i}:=\lambda_{i}(x_{i})$ such that [^1] These are the points in $\lambda$ that intersect the downward leg of the tent function $f_{(x_{1}-y_{1},x_{1}+y_{1})}$. It may happen that the only such point is $(x_{1},y_{1})$, corresponding to the case $n=1$. See Figure 2.
$f_{(x_{1}-y_{1},x_{1}+y_{1})}(x_{i})=\lambda_{i}(x_{i})=y_{i}\quad\mbox{and}\quad D^{-}\lambda_{i}(x_{i})=+1.$
Moreover, let $x_{j}=x_{1}+y_{1}$ for $j\geq n+1$. Define the following sequence of functions $\eta$ as follows
$\eta_{k}(t)=\begin{cases}\lambda_{k+1}(t)&\text{if }t<x_{k+1}\\ \lambda_{k}(t)&\text{if }x_{k+1}\leq t\end{cases}$
for each $k\in\mathbb{N}$. Then:
  1. 1.
    $\eta$ is a landscape sequence.
  2. 2.
    $\eta$ has strictly fewer local maxima than $\lambda$.
  3. 3.
    $\lambda=f_{(x_{1}-y_{1},x_{1}+y_{1})}\oplus\eta$.
Proof.
Item 1. We begin by proving the first assertion of the lemma. It is rather straightforward to verify that $\eta$ satisfies Properties 11,22 and $3$ in Definition 3.11. We will focus on verifying Property $4$.
Let $(t_{\ell},h_{\ell})$ be a critical point in $\eta$, with $h_{\ell}>0$. To prove that the degree at this point is non-negative, we aim to show that each $\eta_{k}$ that attains a local minimum at $(t_{\ell},h_{\ell})$ is uniquely paired with some $\eta_{j}$ that attains a local maximum at $(t_{\ell},h_{\ell})$. We begin with the following observation:
Claim:
If $\eta_{k}$ attains a local minimum at $(t_{\ell},h_{\ell})$, then either $t_{\ell}<x_{k+1}$ or $t_{\ell}>x_{k+1}$; in other words, $t_{\ell}\neq x_{k+1}$.
To prove the claim, we argue by contradiction. Suppose that $t_{\ell}=x_{k+1}$. Then, for $t^{-}$ sufficiently close to $t_{\ell}=x_{k+1}$ with $t^{-}<t_{\ell}$ , we have $\eta_{k}(t^{-})=\lambda_{k+1}(t^{-})$. Since $\eta_{k}$ attains a local minimum at $(t_{\ell},h_{\ell})$, this implies that $\lambda_{k+1}$ must have negative slope near such $t^{-}$.
Now, consider two possibilities: either $x_{k+1}<x_{n+1}$ or $x_{k+1}=x_{n+1}=x_{1}+y_{1}$. If $x_{k+1}<x_{n+1}$, then by definition of $x_{k+1}$, $\lambda_{k+1}$ must have positive slope for $t^{-}<x_{k+1}=t_{\ell}$ sufficiently close to $t_{\ell}$. This contradicts the earlier conclusion that $\lambda_{k+1}$ has negative slope. Hence, we must have $x_{k+1}=x_{n+1}$.
In this case, since $\eta_{k}$ is a local minimum curve and satisfies $\eta_{k}(t^{-})=\lambda_{k+1}(t^{-})$ for $t^{-}<x_{k+1}$, the slope of $\lambda_{k+1}$ near $t^{-}$ must again be negative. But this would imply that $x_{k+1}<x_{n+1}$, since $\lambda_{k+1}$ would then be forced to intersect the tent function $f_{(x_{1}-y_{1},x_{1}+y_{1})}$ at a point earlier than $x_{n+1}$. This contradicts the earlier established fact that $x_{k+1}=x_{n+1}$.
Therefore, both possibilities, $x_{k+1}<x_{n+1}$ or $x_{k+1}=x_{n+1}$, lead to contradictions. We conclude that $t_{\ell}\neq x_{k+1}$; i.e., either $t_{\ell}<x_{k+1}$ or $x_{k+1}<t_{\ell}$. This proves the claim. We now study these two cases. We first show that $\eta_{k}$ is paired with some $\eta_{j}$ that attains a local maximum at $(t_{\ell},h_{\ell})$; we then argue that this pairing is unique.
Case 1:
Suppose $t_{\ell}<x_{k+1}$. Then, $\eta_{k}=\lambda_{k+1}$ in a sufficiently small neighborhood of $t_{\ell}$. This equality implies that $(t_{\ell},h_{\ell})$ is a local minimum point of $\lambda_{k+1}$. Since $\lambda$ has non-negative degree, $\lambda_{k+1}$ can be uniquely paired with a local maximum curve $\lambda_{m}$ at $(t_{\ell},h_{\ell})$ for some $m>k+1$. Moreover, since $t_{\ell}<x_{k+1}\leq x_{m}$, we have $\eta_{m-1}=\lambda_{m}$ in a sufficiently small neighborhood of $t_{\ell}$. Therefore, $\eta_{k}$ is paired with the local maximum curve $\eta_{m-1}$ at $(t_{\ell},h_{\ell})$, in this case.
Case 2:
Suppose that $t_{\ell}>x_{k+1}$. In this case, we can find a small neighborhood of $t_{\ell}$ such that $\eta_{k}=\lambda_{k}$, which implies that $(t_{\ell},h_{\ell})$ is a local minimum of $\lambda_{k}$. Since $\lambda$ has non-negative degree, there exists a unique $m>k$ such that $(t_{\ell},h_{\ell})$ is a local maximum of $\lambda_{m}$. To show that $\lambda_{m}$ corresponds to some $\eta_{j}$, we first observe the following fact: $f_{(x_{1}-y_{1},x_{1}+y_{1})}(t_{\ell})\leq h_{\ell}$. Indeed, if $f_{(x_{1}-y_{1},x_{1}+y_{1})}(t_{\ell})>h_{\ell}$, it follows that
$\lambda_{k+1}(t_{\ell})\geq f_{(x_{1}-y_{1},x_{1}+y_{1})}(t_{\ell})>h_{\ell}=\lambda_{k}(t_{\ell}),$
which is impossible since $\lambda_{k}\geq\lambda_{k+1}$ always. Therefore, $f_{(x_{1}-y_{1},x_{1}+y_{1})}(t_{\ell})\leq h_{\ell}$. This observation implies that $x_{m}\leq t_{\ell}$ (which can also be shown by contradiction similarly). Now, if $x_{m}<t_{\ell}$, then $\eta_{m}=\lambda_{m}$ sufficiently near $t_{\ell}$. Otherwise, if $x_{m}=t_{\ell}$, we have
$\eta_{m-1}(t^{-})$ $=\lambda_{m}(t^{-})\text{ for }t^{-}<t_{\ell}\text{ sufficiently near }t_{\ell}\text{, and }$
$\eta_{m}(t^{+})$ $=\lambda_{m}(t^{+})\text{ for }t^{+}\geq t_{\ell}\text{ sufficiently near }t_{\ell}.$
Since $\eta_{m-1}\geq\eta_{m}$, this forces $\eta_{m-1}(t^{-})=\eta_{m}(t^{-})$ for $t^{-}<t_{\ell}$ near $t_{\ell}$, and thus $\eta_{m}(t)=\lambda_{m}(t)$ for all $t$ in a sufficiently small neighborhood of $t_{\ell}$. Hence, $\eta_{k}$ is paired with $\eta_{m}$ in this case.
Having established that each $\eta_{k}$ pairs with some $\eta_{j}$, we proceed to show that this pairing is unique. Suppose that $\eta_{k}$ and $\eta_{k^{\prime}}$, with $k\neq k^{\prime}$, both attain a local minimum at $(t_{\ell},h_{\ell})$. Using arguments similar to those above, one can show that both $x_{k+1}$ and $x_{k^{\prime}+1}$ must lie on the same side of $t_{\ell}$; that is, either $t_{\ell}<x_{k+1},x_{k^{\prime}+1}$ or $t_{\ell}>x_{k+1},x_{k^{\prime}+1}$. Assuming the former, it is easy to see that $\eta_{k}$ and $\eta_{k^{\prime}}$ are paired with local maximum curves $\eta_{m}$ and $\eta_{m^{\prime}}$ for some distinct $m\neq m^{\prime}$. The latter case follows similarly. In conclusion, we have $\mathrm{deg}_{\eta}(t_{\ell},h_{\ell})\geq 0$, and thus $\eta$ is a landscape sequence.
Item 2.
We claim that $\eta$ admits strictly fewer local maxima than $\lambda$. By construction, every local maximum of $\eta$ is a local maximum of $\lambda$. Moreover, since $\lambda_{1}(t_{1})=h_{1}\neq\eta_{k}(t_{1})$ for any $k$, the point $(t_{1},h_{1})$ is a local maximum of $\lambda$, but not of $\eta$. Therefore, $\eta$ must have strictly fewer local maxima than $\lambda$, and hence at most M1M-1 in total.
Item 3.
We aim to show that $\lambda=f_{(x_{1}-y_{1},x_{1}+y_{1})}\oplus\eta$. For notational simplicity, let us write $f=f_{(x_{1}-y_{1},x_{1}+y_{1})}$ and $\tau=f\oplus\eta$. We claim that $\tau_{k}(t)=\lambda_{k}(t)$ for any $k\in\mathbb{N},t\in\mathbb{R}$. Fix $t\in\mathbb{R}$. There are three cases to consider: either $t<x_{1}$, or $x_{n+1}\leq t$, or $x_{i}\leq t<x_{i+1}$ for some $1\leq i\leq n$. In each case, the key is to determine whether $\tau_{k}(t)$ is realized by the first summand $f$ or the second $\eta$. We present only the last case, as the others follow similarly. So, assume that $x_{i}\leq t<x_{i+1}$ for some $1\leq i\leq n$. Then:
$f(t)$ $=\lambda_{i}(t),$
$\eta_{k}(t)$ $=\lambda_{k}(t)\text{ for }1\leq k\leq i-1,\text{ and }$
$\eta_{k^{\prime}}(t)$ $=\lambda_{k^{\prime}+1}(t)\text{ for }i\leq k^{\prime}.$
This gives the chain of inequalities:
$\eta_{1}(t)=\lambda_{1}(t)\geq\dots\geq\eta_{i-1}(t)=\lambda_{i-1}(t)\geq f(t)=\lambda_{i}(t)\geq\eta_{i}(t)=\lambda_{i+1}(t)\geq\dots$
Hence, $\tau_{k}(t)=\lambda_{k}(t)$ as desired. ∎
With Lemma 3.18 established, we can now prove the Landscape Decomposition Theorem 3.17.
Proof of Theorem 3.17.
The uniqueness claim is straightforward. Indeed, given a landscape sequence of the form $f_{(b_{1},d_{1})}\oplus f_{(b_{2},d_{2})}\oplus\cdots\oplus f_{(b_{N},d_{N})}$, observe that the points $(b_{i},d_{i})$ are exactly the local maxima with strictly positive degree. As the multiset of such points is an intrinsic property of a landscape sequence, any two decompositions can only differ by reordering summands.
It remains to prove the existence of a direct sum decomposition of a landscape sequence $\lambda$. Let $M$ denote the number of local maxima of $\lambda$. We argue by induction on $M$.
To establish the base case, assume $M=1$. Let $(x_{1},y_{1})$ be the unique local maximum of $\lambda$. It must correspond solely to the local maximum of $\lambda_{1}$. Moreover, we must have $\lambda_{1}(x_{1}-y_{1})=0=\lambda_{1}(x_{1}+y_{1})$; otherwise, $\lambda_{1}$ would attain a local minimum in the interval $x_{1}-y_{1},x_{1}+y_{1}]$ which would induce an additional local maximum curve in $\lambda$. Lastly, since there is only one local maximum, we must have $\lambda_{k}=0$ for all $k\geq 2$. This shows that $\lambda=f_{(x_{1}-y_{1},x_{1}+y_{1})}$, proving the base case.
Now assume that every landscape sequence with fewer than M1M-1 local maxima admits a decomposition as in the statement of this theorem, and assume that $\lambda$ has $M$ local maxima. Let $(x_{1},y_{1})$ be the left-most point where $\lambda_{1}$ attains a local maximum. By Lemma 3.18, we have a landscape sequence $\eta$ that admits strictly fewer local maxima than $\lambda$ and satisfies $\lambda=f_{(x_{1}-y_{1},x_{1}+y_{1})}\oplus\eta$. Therefore, by the induction hypothesis, $\eta$ admits a decomposition, and so does $\lambda$. This concludes the proof of existence. ∎
Figure 2: Example demonstrating the effect of removing a tent function from a landscape sequence, as described in Lemma 3.18. Left: Original landscape sequence $\lambda$. The points $(x_{i},y_{i})$ as in the statement of the lemma are annotated. Right: The resulting landscape sequence $\eta$ obtained from $\lambda$ by removing the tent function $f_{(x_{1}-y_{1},x_{1}+y_{1})}$. The colors here correspond to those of the original landscape $\lambda$, changing at the points $x_{i}$, indicating a transition from one landscape curve to the other, in line with the definition of $\eta$, as described in the lemma.
Finally, we leverage the direct sum decomposition to give a proof of Theorem 3.13.
Proof of Theorem 3.13.
Let $Y=\{(b_{i},d_{i})\}_{i=1}^{N}$ be a persistence diagram. It is easy (by definition of $\lambda^{Y}$) to see that $\lambda^{Y}=f_{(b_{1},d_{1})}\oplus f_{(b_{2},d_{2})}\oplus\dots\oplus f_{(b_{N},d_{N})}$. Therefore, $\lambda^{Y}$ is a landscape sequence by Proposition 3.16.
Conversely, let $\lambda$ be a landscape sequence. By Theorem 3.17, there exists a unique multiset of tent functions $\{f_{(b_{1},d_{1})},\dots,f_{(b_{N},d_{N})}\}$ such that
$\lambda=f_{(b_{1},d_{1})}\oplus f_{(b_{2},d_{2})}\oplus\dots\oplus f_{(b_{N},d_{N})}.$
Now, define the persistence diagram $Y=\{(b_{i},d_{i})\}_{i=1}^{N}$ (this is well-defined, by the uniqueness part of the decomposition theorem). By construction, $\lambda^{Y}$ has the same decomposition as $\lambda$. Thus, $\lambda=\lambda^{Y}$. ∎
3.3 Equivalence of Erosion and Landscape Distances
In this section, we prove our main theorem, which equates erosion distance and the landscape distance (5).
Theorem 3.19 (Equivalence of Erosion and Landscape Distances).
The landscape distance between any two persistence landscapes is equal to the erosion distance between their corresponding persistence diagrams. That is, $\|\lambda^{Y}-\lambda^{Y^{\prime}}\|_{\infty}=d_{E}(Y,Y^{\prime})$ for any $Y,Y^{\prime}\in\mathsf{PDgm}$.
The proof follows by realizing the landscape distance as an interleaving distance with respect to a coflow and showing that the resulting structure is equivalent to the one established for persistence diagrams in Section 3.1.2.
Remark 3.20.
In [31, Theorem 4.3], the authors show that the erosion distance can be realized as a Hausdorff distance in the space $\mathbb{R}_{<}^{2}$ of intervals, endowed with $\ell^{\infty}$ distance. After unpacking definitions, one can show that the latter quantity is equal to landscape distance, thereby providing an alternate proof strategy for Theorem 3.19. Our perspective provides a more elementary approach (indeed, [31, Theorem 4.3] is an example of their rather general category-theoretic framework), while leveraging the structural results for persistence landscapes derived above.
3.3.1 Equipping Persistence Landscapes with a Coflow
We now define the poset category of landscape sequences, and equip it with a coflow.
Definition 3.21.
By $\mathsf{Land}=(\mathsf{Land},\leq)$, we denote the poset of landscape sequences $\lambda$ where the poset relation $\leq$ is given by $\lambda\leq\lambda^{\prime}\iff\lambda_{k}(t)\leq\lambda^{\prime}_{k}(t)$ for all $k\in\mathbb{N},t\in\mathbb{R}$.
Let $\varepsilon\geq 0$ and $\lambda=(\lambda_{1},\lambda_{2},\dots)$. Define an operator $T_{\varepsilon}$ on landscape curves by
$T_{\varepsilon}(\lambda_{k}):$ $\mathbb{R}\to\mathbb{R}$ $t\mapsto T_{\varepsilon}(\lambda_{k})(t)=\begin{cases}\lambda_{k}(t)-\varepsilon,&\text{if }\lambda_{k}(t)>\varepsilon\\ 0,&\text{otherwise.}\end{cases}$
That is, $T_{\varepsilon}(\lambda_{k})$ is obtained from $\lambda_{k}$ by shifting the graph of $\lambda_{k}$ downward by $\varepsilon$ units, and truncating as necessary. Observe that if $\lambda_{k}^{\prime}\leq\lambda_{k}$, then $T_{\varepsilon}(\lambda_{k})\leq T_{\varepsilon}(\lambda_{k}^{\prime})$. This operation therefore gives rise to the following well-defined functor.
Definition 3.22.
For each $\varepsilon\geq 0$, define the landscape $\varepsilon$-flow functor $\Lambda_{\varepsilon}:\mathsf{Land}\to\mathsf{Land}$, by $\Lambda_{\varepsilon}\lambda=(T_{\varepsilon}(\lambda_{1}),T_{\varepsilon}(\lambda_{2}),\dots)$.
The proof of the following is straightforward.
Proposition 3.23.
The family $\Lambda=\{\Lambda_{\varepsilon}\}_{\varepsilon\geq 0}$ is a coflow.
The interleaving distance induced by $\Lambda$ is familiar:
Proposition 3.24.
The interleaving distance induced by the coflow $\Lambda$ is equal to the landscape distance. That is,
$d^{\Lambda}_{\mathsf{Land}}(\lambda,\lambda^{\prime})=\|\lambda-\lambda^{\prime}\|_{\infty}$
for all $\lambda,\lambda^{\prime}\in\mathsf{Land}$.
Proof.
Our construction of the interleaving distance $d_{\mathsf{Land}}^{\Lambda}$ as a distance between functions $\mathbb{N}\times\mathbb{R}\to\mathbb{R}$ agrees with the construction used to prove [24, Theorem 3.9], which shows that $L^{\infty}$ distance on $\mathbb{R}^{n}$ can be realized as an interleaving distance. Indeed, small modifications to that proof yield the proof in the setting of this proposition. ∎
3.3.2 Coflow Equivalence of Persistence Landscapes and Persistence Diagrams
Let us denote the landscape map $Y\mapsto\lambda^{Y}$ by $\mathcal{L}:\mathsf{PDgm}\to\mathsf{Land}$. Theorem 3.13 shows that the landscape map is a bijection; let $\mathcal{M}:\mathsf{Land}\to\mathsf{PDgm}$ denote its inverse. In this section, we refine this result and show that $\mathcal{L}$ is a coflow equivalence between the posets $(\mathsf{PDgm},\leq)$ and $(\mathsf{Land},\leq)$.
Proposition 3.25.
Let $Y,Y^{\prime}\in\mathsf{PDgm}$. Then
$Y\leq Y^{\prime}\iff\lambda^{Y}\leq\lambda^{Y^{\prime}}$
That is, morphisms are preserved among the posets $\mathsf{PDgm}$ and $\mathsf{Land}$, so that $\mathcal{L}$ and $\mathcal{M}$ are functors.
Proof.
First, suppose that $Y\leq Y^{\prime}$. Let $k\in\mathbb{N}$ and $t\in\mathbb{R}$, and assume that there exists $h\geq 0$ such that $\mathrm{rank}[Y](t-h,t+h)\geq k$. Then
$k\leq\mathrm{rank}[Y](t-h,t+h)\leq\mathrm{rank}[Y^{\prime}](t-h,t+h).$
Since this holds for all $h$ satisfying the condition, this shows that $\lambda_{k}^{Y}\leq\lambda_{k}^{Y^{\prime}}$; if no such $h$ exists, then the desired conclusion is obvious, as $\lambda_{k}^{Y}=0$. As $k$ was arbitrary, this shows that $\lambda^{Y}\leq\lambda^{Y^{\prime}}$.
Conversely, suppose that $\lambda^{Y}\leq\lambda^{Y^{\prime}}$. Let $I=(b,d)\in\mathbb{R}_{\leq}^{2}$ and suppose that $\mathrm{rank}[Y](I)=k$. If $k=0$, the desired conclusion that $\mathrm{rank}[Y](I)\leq\mathrm{rank}[Y^{\prime}](I)$ is obvious, so let us suppose that $k\geq 1$. Then we have
$k=\mathrm{rank}[Y](b,d)=\mathrm{rank}[Y]\left(\frac{d+b}{2}-\frac{d-b}{2},\frac{d+b}{2}+\frac{d-b}{2}\right).$
This implies
$\frac{d-b}{2}\leq\lambda_{k}^{Y}\left(\frac{d+b}{2}\right)\leq\lambda_{k}^{Y^{\prime}}\left(\frac{d+b}{2}\right),$
which, in turn, implies that $\mathrm{rank}[Y^{\prime}](I)\geq k$. Since $I$ was arbitrary, this implies that $Y\leq Y^{\prime}$. ∎
Proposition 3.26.
The landscape functor $\mathcal{L}:\mathsf{PDgm}\to\mathsf{Land}$ is coflow-equivariant; that is, $\lambda^{\nabla_{\varepsilon}Y}=\Lambda_{\varepsilon}\lambda^{Y}$ for every $Y\in\mathsf{PDgm}$ and $\varepsilon\geq 0$.
Proof.
Let $k\in\mathbb{N}$, $t\in\mathbb{R}$ and $\varepsilon\geq 0$ be arbitrary. We will show that $\lambda_{k}^{\nabla_{\varepsilon}Y}(t)=T_{\varepsilon}(\lambda_{k}^{Y})(t)$. First, assume that $\lambda^{Y}_{k}(t)-\varepsilon>0$. Then
$T_{\varepsilon}(\lambda_{k}^{Y})(t)=$ $\lambda_{k}^{Y}(t)-\varepsilon$
$=$ $\sup\{h\geq 0\,\mid\,\mathrm{rank}[Y](t-h,t+h)\geq k\}-\varepsilon$
$=$ $\sup\{h-\varepsilon\geq 0\,\mid\,\mathrm{rank}[Y](t-h,t+h)\geq k\}$
$=$ $\sup\{h^{\prime}\geq 0\,\mid\,\mathrm{rank}[Y](t-h^{\prime}-\varepsilon,t+h^{\prime}+\varepsilon)\geq k\}$
$=$ $\lambda_{k}^{\nabla_{\varepsilon}Y}(t)$
On the other hand, suppose that $\lambda^{Y}_{k}(t)\leq\varepsilon$. Note that this implies for any $h\geq 0$ such that $\mathrm{rank}[Y](t-h,t+h)\geq k$, we have $h\leq\varepsilon$. Now, let $h\geq 0$ be such that $\mathrm{rank}[\nabla_{\varepsilon}Y](t-h,t+h)\geq k$. We will show that $h$ is forced to be 0. Indeed, by Lemma 3.8, we have
$\mathrm{rank}[Y](t-(h+\varepsilon),t+(h+\varepsilon))\geq k$
but as noted this means $h+\varepsilon\leq\varepsilon$. So, $h=0$. That is, $\lambda^{\nabla_{\varepsilon}Y}_{k}(t)=0=T_{\varepsilon}(\lambda_{k}^{Y})(t)$. ∎
The proof of our main result, Theorem 3.19, now follows easily.
Proof of Theorem 3.19.
Propositions 3.25 and 3.26 show that the landscape map $\mathcal{L}:\mathsf{PDgm}\to\mathsf{Land}$ is a coflow equivariant equivalence (see Definition 2.6). By Proposition 2.7, it follows that the interleaving distance induced by the family $\Lambda$ is equal to the erosion distance; that is, $d_{\mathsf{Land}}^{\Lambda}(\lambda^{Y},\lambda^{Y^{\prime}})=d_{E}(Y,Y^{\prime})$. By Proposition 3.24, $d_{\mathsf{Land}}^{\Lambda}(\lambda^{Y},\lambda^{Y^{\prime}})=\|\lambda^{Y}-\lambda^{Y^{\prime}}\|_{\infty}$, so this completes the proof. ∎
4 Metric Properties of the Erosion Distance
This section collects some topological results regarding the erosion distance, many of which are consequences of our main equivalence result, Theorem 3.19.
Throughout this section, we mildly abuse terminology and refer to the metric on $\mathsf{PDgm}$ which is induced by pulling back the sup norm distance under the landscape map as the landscape distance. We use the notation
$d_{L}(Y,Y^{\prime})\coloneqq\|\lambda^{Y}-\lambda^{Y^{\prime}}\|_{\infty}.$ (6)
This convention is introduced to smoothen certain statements of our results.
Some of our results concern coarse embeddings of metric spaces [27, 42], so we recall the basic definition here. Let $(X,d_{X})$ and $(Y,d_{Y})$ be metric spaces and let $f:X\to Y$ be a map. We say that $f$ is a coarse embedding if there exist non-decreasing maps $\rho^{-},\rho^{+}:[0,\infty)\to[0,\infty)$, with $\lim_{t\to\infty}\rho^{-}(t)=\infty$, such that
$\rho^{-}(d_{X}(x,x^{\prime}))\leq d_{Y}(f(x),f(x^{\prime}))\leq\rho^{+}(d_{X}(x,x^{\prime}))$
for all $x,x^{\prime}\in X$. We say that two metrics $d$ and $d^{\prime}$ on the same set $X$ are coarsely equivalent if the identity map coarsely embeds $d$ into $d^{\prime}$ and vice-versa.
4.1 Topological Equivalence of Persistence Diagrams Under Erosion and Bottleneck Distances
In this subsection, we prove that the bottleneck distance $d_{B}$ (see Section 2.2.1) induces the same topology on persistence diagrams as the erosion distance, $d_{E}$. First, note that the bottleneck distance serves as an upper bound for the erosion distance, i.e., one has $d_{E}(Y,Y^{\prime})\leq d_{B}(Y,Y^{\prime})$ for any pair of persistence diagrams $Y$ and $Y^{\prime}$ (see [46, Lemma 6.4]). This implies that a bottleneck-distance open ball, $B_{d_{B}}(Y,r)$, lies inside an erosion-distance open ball, $B_{d_{E}}(Y,r)$. In other words, every $d_{E}$-open set $U$ is a $d_{B}$-open set. However, the bottleneck and erosion distances are not bi-Lipschitz equivalent. In fact, we can say something stronger:
Proposition 4.1.
The bottleneck and erosion distances are not coarsely equivalent.
Proof.
We will show that there is no function $\rho^{-}:[0,\infty)\to[0,\infty)$ with $\lim_{t\to\infty}\rho^{-}(t)=\infty$ such that
$\rho^{-}(d_{B}(Y,Y^{\prime}))\leq d_{E}(Y,Y^{\prime})$
for all $Y,Y^{\prime}\in\mathsf{PDgm}$.
In the proof of [4, Proposition 6.1], Bubenik constructs, for any positive integer $n$, a pair of diagrams $Y_{n}$ and $Y_{n}^{\prime}$ such that
$\|\lambda^{Y_{n}}-\lambda^{Y_{n}^{\prime}}\|_{\infty}=1\quad\mbox{and}\quad d_{B}(Y_{n},Y_{n}^{\prime})=2n+1.$
By Theorem 4.2, the first equality implies $d_{E}(Y_{n},Y_{n}^{\prime})=1$. Now let $\rho^{-}:[0,\infty)\to[0,\infty)$ be any function with $\lim_{t\to\infty}\rho^{-}(t)=\infty$. Choose $n$ such that $\rho^{-}(2n+1)>1$. Then
$d_{E}(Y_{n},Y_{n}^{\prime})=1<\rho^{-}(d_{B}(Y_{n},Y_{n}^{\prime})).$
This proves our initial non-existence claim. ∎
Despite the fact that the bottleneck and erosion distances are not coarsely equivalent, we now prove that they are locally equivalent.
Theorem 4.2 (Local Isometry of Bottleneck and Erosion Distances).
The bottleneck and erosion distances on $\mathsf{PDgm}$ are locally isometric. In particular, let $Y=\{(b_{i},d_{i})\}_{i=1}^{N}$ be a persistence diagram and set
$r_{Y}\coloneqq\frac{1}{2}\min\left\{\lvert b_{i}-b_{j}\rvert,\lvert d_{i}-d_{j}\rvert,\frac{d_{i}-b_{i}}{2}\,:b_{i}\neq b_{j},d_{i}\neq d_{j}\right\}.$
If $Y^{\prime}\in\mathsf{PDgm}$ satisfies $d_{B}(Y,Y^{\prime})<r_{Y}$ or $d_{E}(Y,Y^{\prime})<r_{Y}$ then $d_{B}(Y,Y^{\prime})=d_{E}(Y,Y^{\prime})$.
Terminology 4.3.
We introduce some terminology related to the definition of bottleneck distance, given in Section 2.2.1.
  • We will refer to a triplet $(S,S^{\prime},\phi)$, with $S\subset Y$, $S^{\prime}\subset Y^{\prime}$, and $\phi:S\to S^{\prime}$ a bijection, as a partial matching between $Y$ and $Y^{\prime}$.
  • Given a partial matching one can interpret the quantity
    $\max\left\{\max_{(b,d)\in S}d_{\infty}((b,d),\phi(b,d)),\;\max_{(b,d)\notin S}\frac{(d-b)}{2},\;\max_{(b^{\prime},d^{\prime})\notin S^{\prime}}\frac{(d^{\prime}-b^{\prime})}{2}\right\}$
    as the cost of $(S,S^{\prime},\phi)$ and denote it as $\mathrm{cost}(\phi)$.
  • We say that a partial matching $(S,S^{\prime},\phi)$ is optimal if it achieves the minimum cost.
  • By $(b,d)\overset{\phi}{\longleftrightarrow}(b^{\prime},d^{\prime})$, we will denote that $\phi(b,d)=(b^{\prime},d^{\prime})$ and say that $(b,d)$ and $(b^{\prime},d^{\prime})$ are matched under $\phi$. Similarly, by $(b,d)\overset{\phi}{\longleftrightarrow}\Delta$, will denote that $(b,d)$ is matched with the diagonal $\Delta$ under $\phi$.
By choosing $r$ sufficiently small, we guarantee a minimum amount of distance among the birth coordinates; similarly, the death coordinates are also separated by a certain distance. Formally, we state this observation as a lemma below. Its proof is immediate.
Lemma 4.4.
Let $Y=\{(b_{i},d_{i})\}_{i=1}^{N}\in\mathsf{PDgm}$ and $r_{Y}$ be given as in Theorem 4.2. Let $(b_{i},d_{i})\neq(b_{j},d_{j})$ be two distinct birth-death pairs.
  1. 1.
    If $b_{i}\neq b_{j}$, then $2r_{Y}\leq|b_{i}-b_{j}|$
  2. 2.
    Similarly, if $d_{i}\neq d_{j}$, then $2r_{Y}\leq|d_{i}-d_{j}|$
To prove Theorem 4.2, we rely on the following lemma which characterizes open balls with sufficiently small radius in the space of persistence diagrams equipped with either $d_{E}$ or $d_{B}$. For an illustration of such an open ball, we refer the reader to Figure 3.
Lemma 4.5 (Open Ball Lemma).
Let $Y=\{(b_{i},d_{i})\}_{i=1}^{N}\in\mathsf{PDgm}$ and let $r>0$. We let $U(Y,r)$ denote the set of $Y^{\prime}\in\mathsf{PDgm}$ satisfying the following two conditions:
  1. 1.
    For each $(b,d)\in Y$ with multiplicity $m$, there are exactly $m$ birth-death pairs in $Y^{\prime}$ (counted with multiplicity) that lie in the $r$-unit box
    $Q\coloneqq\{(x,y)\in\mathbb{R}^{2}_{<}\,\mid\,d_{\infty}((b,d),(x,y))<r\}$
    centered at $(b,d)$.
  2. 2.
    Each $(b^{\prime},d^{\prime})\in Y^{\prime}$ lies either in one of the open boxes
    $Q_{i}\coloneqq\{(x,y)\in\mathbb{R}^{2}_{<}\,\mid\,d_{\infty}((b_{i},d_{i}),(x,y))<r\}$
    for some $1\leq i\leq N$, or in the open band
    $B=\{(x,y)\in\mathbb{R}^{2}_{<}\,\mid\frac{y-x}{2}<r\}$
    above the diagonal.
Let $r_{Y}$ be as in the statement of Theorem 4.2. If $r\leq r_{Y}$ then the open balls with respect to bottleneck and erosion distances satisfy
$B_{d_{B}}(Y,r)=B_{d_{E}}(Y,r)=U(Y,r).$
In particular, $d_{B}$ and $d_{E}$ are topologically equivalent.
Proof.
Let us first show that $B_{d_{B}}(Y,r)=U(Y,r)$. If $Y^{\prime}\in B_{d_{B}}(Y,r)$, then there exists a matching $\phi$ of $Y$ and $Y^{\prime}$ with cost less than $r<r_{Y}$. It is not hard to show that this forces $Y^{\prime}$ to satisfy the two properties defining $U(Y,r)$, hence $Y^{\prime}\in U(Y,r)$. Conversely, if $Y^{\prime}\in U(Y,r)$, then one can construct a matching $\phi$ with cost less than $r$, hence $d_{B}(Y,Y^{\prime})<r$.
Next we show that $B_{d_{E}}(Y,r)=U(Y,r)$. The easy direction begins with the assumption that $Y^{\prime}\in U(Y,r)$. Then, by the above, $Y^{\prime}\in B_{d_{B}}(Y,r)$, so $d_{E}(Y,Y^{\prime})\leq d_{B}(Y,Y^{\prime})<r$, and $Y^{\prime}\in B_{d_{E}}(Y,r)$. The remainder of the proof focuses on the converse, which takes significantly more work.
Suppose that $d_{E}(Y,Y^{\prime})<r$. Then there exists $\varepsilon\geq 0$ with $d_{E}(Y,Y^{\prime})\leq\varepsilon<r$ such that the two rank conditions in Equation 4 are satisfied. Let $(b,d)\in Y$ be an arbitrary pair with multiplicity $m\geq 1$. We consider the four distinct corners of the open box $Q$; that is
  • top-left: $I_{1}=(b-r,d+r)$,
  • top-right: $I_{2}=(b+r,d+r)$,
  • bottom-left: $I_{3}=(b-r,d-r)$,
  • bottom-right: $I_{4}=(b+r,d-r)$.
These give rise to four disjoint regions below:
  • Region $1$: $\mathsf{Green}\coloneqq\{(x,y)\in\mathbb{R}^{2}_{<}\,\mid\,(x,y)\text{ properly contains }I_{1}\}$
  • Region $2$: $\mathsf{Yellow}\coloneqq\{(x,y)\in\mathbb{R}^{2}_{<}\,\mid\,(x,y)\text{ properly contains }I_{3},\text{ but not $I_{1}$}\}$
  • Region $3$: $\mathsf{Orange}\coloneqq\{(x,y)\in\mathbb{R}^{2}_{<}\,\mid\,(x,y)\text{ properly contains }I_{2},\text{ but not }I_{1}\}$
  • Region $4$: the $r$-unit open box $Q$.
An example scenario is shown in Figure 3. Assume that $Y$ admits exactly $n_{1}\geq 0$ pairs in $\mathsf{Green}$, $n_{2}\geq 0$ pairs in $\mathsf{Yellow}$ and $n_{3}\geq 0$ pairs in $\mathsf{Orange}$. Since $(b,d)$ has multiplicity $m$, and it properly contains $I_{4}$, we have $\mathrm{rank}[Y](I_{4})=m+n_{1}+n_{2}+n_{3}$. Define $\mathrm{Shrink}_{\varepsilon}(x,y)\coloneqq(x+\varepsilon,y-\varepsilon)$. Since $r$ is small enough and $\varepsilon<r$, Lemma 4.4 implies that
$\mathrm{rank}[\nabla_{\varepsilon}Y](I)=\mathrm{rank}[Y](I)=\mathrm{rank}[Y](\mathrm{Shrink}_{\varepsilon}I)$
and moreover $\mathrm{Grow}_{\varepsilon}\circ\mathrm{Shrink}_{\varepsilon}(I)=I$ for $I\in\{I_{1},I_{2},I_{3},I_{4}\}$. Since $\varepsilon$ satisfies the rank conditions, we have
$\mathrm{rank}[\nabla_{\varepsilon}Y](I)\leq\mathrm{rank}[Y^{\prime}](I)\leq\mathrm{rank}[Y](\mathrm{Shrink}_{\varepsilon}(I))$
which in turn yields the equality
$\mathrm{rank}[Y^{\prime}](I)=\mathrm{rank}[Y](I)$
for all $I\in\{I_{1},I_{2},I_{3},I_{4}\}$. Explicitly, this means:
  • $\mathrm{rank}[Y^{\prime}](I_{1})=\mathrm{rank}[Y](I_{1})=n_{1}$; that is $Y^{\prime}$ must admit exactly $n_{1}$ pairs that properly contain $I_{1}$, all of which lie in $\mathsf{Green}$ by definition.
  • $\mathrm{rank}[Y^{\prime}](I_{2})=\mathrm{rank}[Y](I_{2})=n_{1}+n_{3}$; that is $Y^{\prime}$ must admit exactly $n_{1}+n_{3}$ pairs that properly contain $I_{2}$. This implies that there are exactly $n_{3}$ pairs in $Y^{\prime}$ that lie in $\mathsf{Orange}$.
  • $\mathrm{rank}[Y^{\prime}](I_{3})=\mathrm{rank}[Y](I_{3})=n_{1}+n_{2}$; that is $Y^{\prime}$ must admit exactly $n_{1}+n_{2}$ pairs that properly contain $I_{3}$. This implies that there are exactly $n_{2}$ pairs in $Y^{\prime}$ that lie in $\mathsf{Yellow}$.
  • $\mathrm{rank}[Y^{\prime}](I_{4})=\mathrm{rank}[Y](I_{4})=m+n_{1}+n_{2}+n_{3}$; that is $Y^{\prime}$ must admit exactly $m+n_{1}+n_{2}+n_{3}$ pairs that properly contain $I_{4}$.
In particular, this implies that $Y^{\prime}$ admits exactly $m$ pairs in $Q$, proving the first assertion.
To prove the second assertion, we argue by contradiction. Suppose that $Y^{\prime}$ admits a pair $(b^{\prime},d^{\prime})$ such that $(b^{\prime},d^{\prime})\notin Q_{i}$ for any $i$, and $(b^{\prime},d^{\prime})\notin B$. Consider
$I=\left(\frac{b^{\prime}+d^{\prime}}{2},\frac{b^{\prime}+d^{\prime}}{2}\right)\in\mathbb{R}_{\leq}^{2}.$
Assume that there are exactly $m\geq 0$ pairs in $Y$ that properly contain $I$; that is $\mathrm{rank}[Y](I)=m$. By assumption, each such pair lies at distance at least $r$ from $(b^{\prime},d^{\prime})$. Hence, by the first assertion of this lemma, we can find at least $m$ pairs in $Y^{\prime}$ (distinct from $(b^{\prime},d^{\prime})$) that properly contain $I$. Moreover, since $\mathrm{Grow}_{\varepsilon}(I)$ lies in the open band $B$ and $Q_{1},\dots,Q_{n},B$ are pairwise disjoint, these pairs in $Y^{\prime}$ also properly contain $\mathrm{Grow}_{\varepsilon}(I)$. We then obtain the contradiction
$m+1\leq\mathrm{rank}[\nabla_{\varepsilon}Y^{\prime}](I)\leq\mathrm{rank}[Y](I)=m.$
Thus, $Y^{\prime}$ cannot admit such a pair $(b^{\prime},d^{\prime})$. This completes the proof of the lemma. ∎
Proof of Theorem 4.2.
First suppose that $d_{B}(Y,Y^{\prime})<r_{Y}$. If $d_{B}(Y,Y^{\prime})=0$, then the statement is trivial, as $d_{E}(Y,Y^{\prime})\leq d_{B}(Y,Y^{\prime})$ holds, in general. So, assume $d_{B}(Y,Y^{\prime})\neq 0$. We aim to show that for any $0\leq\varepsilon<d_{B}(Y,Y^{\prime})$, one of the rank inequalities in the definition of erosion distance fails; that is, there exists some $I\in\mathbb{R}^{2}_{\leq}$ for which
[…]
This will imply the desired equality $d_{E}(Y,Y^{\prime})=d_{B}(Y,Y^{\prime})$.
Let $\phi$ be an optimal matching between $Y$ and $Y^{\prime}$, i.e., $\mathrm{cost}(\phi)=d_{B}(Y,Y^{\prime})$. Without loss of generality, suppose this cost is attained at a matched pair $(b,d)\overset{\phi}{\longleftrightarrow}(b^{\prime},d^{\prime})$ with
$\mathrm{cost}(\phi)=\max\{|b-b^{\prime}|,|d-d^{\prime}|\}=|b-b^{\prime}|=b^{\prime}-b$
The other cases—such as the cost attained by $|d-d^{\prime}|$ or by a diagonal match in $Y^{\prime}$—follow by similar arguments as below.
Set $I=(b+\varepsilon,d-r_{Y})$, which by definition of $r_{Y}$, lies in $\mathbb{R}^{2}_{\leq}$. Recalling Equation 1 and Terminology 2.1, notice that $(b,d)\in Y$ properly contains $\mathrm{Grow}_{\varepsilon}(I)$ i.e. $(b,d)$ contains $\mathrm{Grow}_{\varepsilon}(I)$ as an interval, with $d-r_{Y}+\varepsilon<d$. In contrast, $(b^{\prime},d^{\prime})$ does not properly contain II because $b^{\prime}-b>\varepsilon$. Furthermore:
Claim:
Each birth-death pair $(b^{\prime}_{i},d^{\prime}_{i})\in Y^{\prime}$ that properly contains $I$ is uniquely matched with a birth-death pair $(b_{i},d_{i})\in Y$ that properly contains $\mathrm{Grow}_{\varepsilon}(I)$. Moreover, $(b_{i},d_{i})$ differs from $(b,d)$ either because it is a distinct point in $\mathbb{R}^{2}_{<}$, or because it is a separate copy of the same point $(b,d)$, due to multiplicity in the birth-death pairs.
To prove this claim, let $(b^{\prime}_{i},d^{\prime}_{i})\in Y^{\prime}$ be such a birth-death pair. We first observe that
$d^{\prime}_{i}-b^{\prime}_{i}>d-r_{Y}-b^{\prime}_{i}\geq(d-b)-r_{Y}-\varepsilon\geq 4r_{Y}-r_{Y}-\varepsilon>2r_{Y}$
where the first two inequalities follow from the assumption that $(b^{\prime}_{i},d^{\prime}_{i})$ properly contains $I$, and the latter two follow from the definition of $r_{Y}$ and the assumption that $\varepsilon<d_{B}(Y,Y^{\prime})<r_{Y}$, respectively. This implies that the cost of matching $(b^{\prime}_{i},d^{\prime}_{i})$ to the diagonal would exceed $r_{Y}$. Therefore, it must instead be matched with a unique birth-death pair $(b_{i},d_{i})\in Y$; that is $(b^{\prime}_{i},d^{\prime}_{i})\overset{\phi}{\longleftrightarrow}(b_{i},d_{i})$. Note that since $(b^{\prime},d^{\prime})$ does not properly contain $I$ and $(b,d)\overset{\phi}{\longleftrightarrow}(b^{\prime},d^{\prime})$, it follows that $(b_{i},d_{i})$ differs from $(b,d)$ in the sense explained above. We now establish that $(b_{i},d_{i})$ properly contains $\mathrm{Grow}_{\varepsilon}(I)$; that is $b_{i}\leq b$ and $d-r_{Y}+\varepsilon<d_{i}$. We prove this by way of contradiction as follows:
  • Suppose $b_{i}>b$. Then, by Lemma 4.4, along with the conditions $r_{Y}>\varepsilon$ and $(b^{\prime}_{i},d^{\prime}_{i})$ properly containing $I$, we have
    $b_{i}\geq b+2r_{Y}=b+r_{Y}+r_{Y}>b+\varepsilon+r_{Y}\geq b^{\prime}_{i}+r_{Y}$
    which implies that $|b_{i}-b^{\prime}_{i}|>r_{Y}$. Since $(b^{\prime}_{i},d^{\prime}_{i})\overset{\phi}{\longleftrightarrow}(b_{i},d_{i})$, this gives $\mathrm{cost}(\phi)>r_{Y}$. However, this is impossible because $\mathrm{cost}(\phi)<r_{Y}$. Therefore, we must have $b_{i}\leq b$.
  • Suppose now that $d_{i}\leq d-r_{Y}+\varepsilon$. This means that $d_{i}<d$ since $\varepsilon<r_{Y}$. By similar arguments as above, we obtain:
    $d_{i}\leq d-2r_{Y}=d-r_{Y}-r_{Y}<d^{\prime}_{i}-r_{Y}$
    thus
    $r_{Y}<d^{\prime}_{i}-d_{i}=|d^{\prime}_{i}-d_{i}|$
    but since $(b^{\prime}_{i},d^{\prime}_{i})\overset{\phi}{\longleftrightarrow}(b_{i},d_{i})$, this cannot occur. We must have $d-r_{Y}+\varepsilon<d_{i}$.
This proves the claim. Now assume that $\mathrm{rank}[Y^{\prime}](I)=m$; that is, there are exactly $m$ birth-death pairs in $Y^{\prime}$ that properly contain $I$. By the claim above, this implies that there are at least $m$ birth-death pairs in $Y$, distinct from $(b,d)$, that properly contain $\mathrm{Grow}_{\varepsilon}(I)$. Hence, we have
$\mathrm{rank}[\nabla_{\varepsilon}Y](I)=\mathrm{rank}[Y](\mathrm{Grow}_{\varepsilon}(I))\geq m+1>m=\mathrm{rank}[Y^{\prime}](I)$
Therefore, the rank inequality condition of the erosion distance definition is violated for $\varepsilon<d_{B}(Y,Y^{\prime})$, which forces equality of the distances since we also have $d_{E}(Y,Y^{\prime})\leq d_{B}(Y,Y^{\prime})$.
To complete the proof, now assume that $d_{E}(Y,Y^{\prime})<r_{Y}$. Then $d_{B}(Y,Y^{\prime})<r_{Y}$, by Lemma 4.5. By the proof above, it follows that $d_{E}(Y,Y^{\prime})=d_{B}(Y,Y^{\prime})$ as desired. ∎
Figure 3: An example pair of persistence diagrams $Y$ and $Y^{\prime}$ such that $d_{E}(Y,Y^{\prime})<r_{Y}$ where $r_{Y}$ is as in Theorem 4.2. Birth-death pairs in $Y$ are shown as blue points in $\mathbb{R}^{2}_{<}$, and those in $Y^{\prime}$ are shown in red. The regions $\mathsf{Green},\mathsf{Yellow},\mathsf{Orange}$ and $Q$, together with the points $I_{1},I_{2},I_{3},I_{4}\in\mathbb{R}^{2}_{<}$, and the birth-death pair $(b,d)$ are as defined in the proof of Lemma 4.5. We set $r=r_{Y}$ for this illustration. The $r_{Y}$-unit open boxes surrounding each pair in $Y$ together with the open band $B$ (shown in gray above the diagonal $\Delta$) form the open ball with radius $r$ centered at $Y$ with respect to either metric $d_{E}$ or $d_{B}$. For each pair $(b_{i},d_{i})$ in $Y$, one has exactly as many pairs in $Y^{\prime}$ as the multiplicity of $(b_{i},d_{i})$ in the corresponding box. In addition, $Y^{\prime}$ may have any number of birth-death pairs in the open band $B$.
4.2 Erosion Distance is not a Length Metric
As a corollary to Theorem 4.2, we prove that the space of persistence diagrams with the erosion distance, $(\mathsf{PDgm},d_{E})$, is not a length space, and, in fact, its intrinsic metric is the bottleneck distance. The following recalls some standard definitions from metric geometry (see, e.g., [9, 3]).
Definition 4.6.
Let $(X,d)$ be a metric space. Denote by $\mathrm{len}_{d}(\alpha)$ the length of a continuous path $\alpha:[0,1]\to X$. That is,
$\mathrm{len}_{d}(\alpha)=\sup\bigg{\{}\sum d\left(\alpha(t_{i}),\alpha(t_{i+1})\right)\,\mid\,0=t_{0}<t_{1}<\cdots<t_{n+1}=1\bigg{\}}.$
The intrinsic metric associated to $d$ is defined by
$\hat{d}(x,x^{\prime})\coloneqq\inf\{\mathrm{len}_{d}(\alpha)\mid\alpha(0)=x,\;\alpha(1)=x^{\prime}\},$ (8)
where the infimum is over all continuous paths into $X$ satisfying the given boundary conditions. If $\hat{d}=d$, then $(X,d)$ is called a length space. If $(X,d)$ is a length space such that the infimum in (8) is always realized, then $(X,d)$ is called a geodesic space. In this case, an infimizing path $\alpha$ is called a geodesic, and it satisfies
$d(\alpha(s),\alpha(t))=|s-t|d(x,x^{\prime})$
for all $s,t\in[0,1]$.
Theorem 4.7.
The intrinsic metric corresponding to erosion distance $\hat{d}_{E}$ is bottleneck distance $d_{B}$. That is,
$\hat{d}_{E}(Y,Y^{\prime})=d_{B}(Y,Y^{\prime})$
for all $Y,Y^{\prime}\in\mathsf{PDgm}$.
Proof.
Let $Y,Y^{\prime}\in\mathsf{PDgm}$ be two persistence diagrams. It suffices to show that
$\mathrm{len}_{d_{E}}(\alpha)=\mathrm{len}_{d_{B}}(\alpha)$
for any rectifiable path $\alpha:[0,1]\to\mathsf{PDgm}$ from $Y$ to $Y^{\prime}$. This gives the desired result:
$\hat{d}_{E}(Y,Y^{\prime})=\inf\mathrm{len}_{d_{E}}(\alpha)=\inf\mathrm{len}_{d_{B}}(\alpha)=d_{B}(Y,Y^{\prime})$
where the last equality follows from the fact that $(\mathsf{PDgm},d_{B})$ is geodesic (see, e.g., [17, 16]).
Let us prove our claim that $\mathrm{len}_{d_{E}}(\alpha)=\mathrm{len}_{d_{B}}(\alpha)$. Let $\alpha$ be a rectifiable path from $Y$ to $Y^{\prime}$, and denote $\alpha(t)=Y_{t}$ for $t\in[0,1]$, with $Y_{0}=Y$ and $Y_{1}=Y^{\prime}$. First, note that we will refer to a partition
$P=\{s_{0},s_{1},\dots,s_{m+1}\}$
of the interval $[0,1]$ as a good partition if $d_{E}$ and $d_{B}$ satisfy
$d_{E}(\alpha(s_{i}),\alpha(s_{i+1}))=d_{B}(\alpha(s_{i}),\alpha(s_{i+1}))$
for all $0\leq i\leq m$.
By Theorem 4.2, for any point $t\in[0,1]$, there exists a neighborhood of $\alpha(t):=Y_{t}\in\mathsf{PDgm}$ in which $d_{E}$ and $d_{B}$ measure distances to $Y_{t}$ the same. That is
$d_{E}(Y_{t},Y^{\prime\prime})=d_{B}(Y_{t},Y^{\prime\prime})$
for any $Y^{\prime\prime}$ in that neighborhood. These neighborhoods form an open cover of $\alpha([0,1])$ which is a compact subset of $\mathsf{PDgm}$. Therefore, we can extract a finite subcover. Moreover, we can find points
$0=t_{0}<t_{1}<\cdots<t_{n+1}=1$
such that $t_{i}$ and $t_{i+1}$ lie in the same neighborhood for each $i$. This implies that
$P:=\{t_{0},t_{1},\dots,t_{n+1}\}$
forms a good partition. Therefore
$\sum_{i=1}^{n}d_{E}(\alpha(t_{i}),\alpha(t_{i+1}))=\sum_{i=1}^{n}d_{B}(\alpha(t_{i}),\alpha(t_{i+1}))$
By similar arguments as above, for any arbitrary partition $P^{\prime}=\{t^{\prime}_{0},t^{\prime}_{1},\dots,t^{\prime}_{k+1}\}$ we can find a corresponding good partition $\hat{P^{\prime}}=\{s_{0},s_{1},\dots,s_{m+1}\}$ simply by taking the union of points in $P$ with $P^{\prime}$ so that
$\sum_{i=1}^{m}d_{B}\left(\alpha(s_{i}),\alpha(s_{i+1})\right)=\sum_{i=1}^{m}d_{E}\left(\alpha(s_{i}),\alpha(s_{i+1})\right)\geq\sum_{i=1}^{k}d_{E}\left(\alpha(t^{\prime}_{i}),\alpha(t^{\prime}_{i+1})\right)=\sum_{i=1}^{k}d_{B}\left(\alpha(t^{\prime}_{i}),\alpha(t^{\prime}_{i+1})\right)$
In other words, when computing the lengths $\mathrm{len}_{d_{E}}(\alpha)$ and $\mathrm{len}_{d_{B}}(\alpha)$ , we can instead take the supremum over the set of good partitions. This implies that
$\mathrm{len}_{d_{E}}(\alpha)=\mathrm{len}_{d_{B}}(\alpha)$
as desired. ∎
We have two immediate corollaries.
Corollary 4.8.
The space $(\mathsf{PDgm},d_{E})$ is not a length space.
Proof.
As was observed in Section 2.2.3, $d_{E}\neq d_{B}$, in general. Since the intrinsic metric of $d_{E}$ is not itself, $(\mathsf{PDgm},d_{E})$ is not a length space. ∎
Applying Theorem 4.2 yields the following reinterpretation.
Corollary 4.9.
The space $(\mathsf{PDgm},d_{L})$ is not a length space. In fact, the bottleneck distance is the intrinsic metric associated to the landscape distance.
4.3 Birth-Zero Diagrams
We study a specific subset of persistence diagrams, defined as follows.
Definition 4.10.
By $\mathsf{PDgm}^{0}$, we denote the proper subset of persistence diagrams of the form
$Y=\{(0,d_{i})\}_{i=1}^{N}$
with $d_{1}\geq d_{2}\geq\cdots\geq d_{N}>0$. We refer to such diagrams as birth-zero persistence diagrams.
Remark 4.11.
When one works with persistence diagrams that arise from the Vietoris-Rips filtration in homology degree zero $H_{0}$, the resulting diagram is a birth-zero persistence diagram. For this reason, it is sometimes referred to as a zero-degree persistence diagram, as in [39, Patrangenaru et al., 2019].
On this subspace of persistence diagrams, we are able to derive a stronger connection between the bottleneck and erosion distances, and to relate them to an alternative vectorization introduced by Patrangenaru et al.  in [39]. These results provide key tools for addressing coarse non-embeddability questions in Section 4.4 below.
Although Proposition 4.1 states that bottleneck and erosion distances are not equivalent (even in a weak sense) on the full space $\mathsf{PDgm}$, the next result shows that they agree on the subspace $\mathsf{PDgm^{0}}$.
Theorem 4.12.
The bottleneck distance and the erosion distance agree on $\mathsf{PDgm}^{0}$; that is, for any $Y,Y^{\prime}\in\mathsf{PDgm}^{0}$, we have
$d_{B}(Y,Y^{\prime})=d_{E}(Y,Y^{\prime}).$
Proof.
First consider diagrams consisting of single points $\{(0,d)\}$ and $\{(0,d^{\prime})\}$. It is not hard to show, by direct computation, that
$d_{L}(\{(0,d)\},\{(0,d^{\prime})\})=\|f_{(0,d)}-f_{(0,d^{\prime})}\|_{\infty}=\min\{|d-d^{\prime}|,\max\{d/2,d^{\prime}/2\}\}.$
On the other hand, this is also the formula for $d_{B}(\{(0,d)\},\{(0,d^{\prime})\})$ (there are only two choices of matchings for these diagrams, whose costs are the elements of the set given above).
Now consider the general case $Y=\{(0,d_{i})\}_{i=1}^{N}$ and $Y^{\prime}=\{(0,d_{i}^{\prime})\}_{i=1}^{N^{\prime}}$. Without loss of generality, assume that $N^{\prime}\geq N$ and set $d_{N+1}=\cdots=d_{N^{\prime}}=0$. Then, by the fact observed above and Theorem 3.19,
$\begin{split}d_{E}(Y,Y^{\prime})=d_{L}(Y,Y^{\prime})&=\max_{i=1,\ldots,N^{\prime}}d_{L}(\{(0,d_{i})\},\{(0,d_{i}^{\prime})\})\\ &=\max_{i=1,\ldots,N^{\prime}}d_{B}(\{(0,d_{i})\},\{(0,d_{i}^{\prime})\})\geq d_{B}(Y,Y^{\prime}),\end{split}$
where the last inequality follows because the quantity on its left hand side is the cost of a particular matching. Indeed, it is easy to verify that
$\mathrm{cost}(\phi)=\max_{i=1,\dots,N^{\prime}}\min\{|d_{i}-d^{\prime}_{i}|,\max\{d_{i}/2,d^{\prime}_{i}/2\}\}$
where the matching $\phi$ pairs $(0,d_{i})\overset{\phi}{\longleftrightarrow}(0,d^{\prime}_{i})$ if and only if
$|d_{i}-d^{\prime}_{i}|\leq\max\{d_{i}/2,d^{\prime}_{i}/2\}$
and otherwise both points are matched with the diagonal, i.e.,
[…]
for each $i=1,\dots,N$. Moreover, for indices $i>N$, the points $(0,d^{\prime}_{i})$ are matched with $\Delta$. Since $d_{E}\leq d_{B}$, in general, this completes the proof. ∎
Remark 4.13.
A result similar to Theorem 4.12, comparing bottleneck distance and a certain interleaving distance between Betti curves on the subspace of birth-zero persistence diagrams, is obtained in [30, Theorem 4.14]. The proof strategies are similar, but comparing the explicit calculations of [30, Example 4.15] shows that their result is distinct from ours.
Remark 4.14.
The computational complexity of the bottleneck distance was initially explored by Efrat et al.  [26], and more recently Kerber et al. [29, Theorem 3.1] proposed an efficient algorithm. While we do not delve into a detailed discussion of the bottleneck distance’s computational complexity here, it is worth noting that the proof of Theorem 4.12 leads to a simple formulation—which suggest a linear-time computation—of the distances $d_{B}=d_{E}=d_{L}$ on $\mathsf{PDgm}^{0}$, as we record in the following corollary. Indeed, the partial matching constructed in the proof of Theorem 4.12 is optimal.
Corollary 4.15.
Let $Y=\{(0,d_{i})\}_{i=1}^{N}$ and $Y^{\prime}=\{(0,d_{i}^{\prime})\}_{i=1}^{N^{\prime}}$, and without loss of generality, assume that $N^{\prime}\geq N$. Set $d_{N+1}=\cdots=d_{N^{\prime}}=0$. For $d_{\bullet}=d_{B}=d_{E}=d_{L}$, we have
$d_{\bullet}(Y,Y^{\prime})=\max_{i=1,\dots,N^{\prime}}\min\bigg{\{}|d_{i}-d^{\prime}_{i}|,\max\{d_{i}/2,d^{\prime}_{i}/2\}\bigg{\}}.$
Birth-zero persistence diagrams can be represented as sequences in $\ell^{\infty}$ by listing their death coordinates followed by infinitely many zeros. This is known as the death-vectorization embedding due to Patrangenaru et. al. [39]. An easy corollary of Theorem 4.12 is that this embedding is bi-Lipschitz. Let us now formalize this result.
Definition 4.16.
[39] The death-vectorization map $\mathrm{DV}:\mathsf{PDgm}^{0}\to\ell^{\infty}$, is defined by
$\mathrm{DV}\left(\left\{(0,d_{i})\right\}_{i=1}^{n}\right)=(d_{1},\dots,d_{n},0,\dots).$
Remark 4.17.
Let $\ell^{\infty}_{\geq}\subset\ell^{\infty}$ denote the subset of sequences of the form $(a_{1},\dots,a_{n},0,\dots)$ with $a_{1}\geq\dots\geq a_{n}>0$ for some $n\in\mathbb{N}$. Notice that this is the image of the death-vectorization map, and that $\mathrm{DV}:\mathsf{PDgm}^{0}\to\ell^{\infty}_{\geq}$ is a bijection.
Corollary 4.18.
The death-vectorization map satisfies
$d_{\bullet}(Y,Y^{\prime})\leq\|\mathrm{DV}(Y)-\mathrm{DV}(Y^{\prime})\|_{\infty}\leq 2\,d_{\bullet}(Y,Y^{\prime})$
for all $Y,Y^{\prime}\in\mathsf{PDgm}^{0}$, where $d_{\bullet}=d_{L}=d_{B}=d_{E}$.
Proof.
Let $Y=\{(0,d_{i})\}_{i=1}^{N}$ and $Y^{\prime}=\{(0,d_{i}^{\prime})\}_{i=1}^{N^{\prime}}$ be an arbitrary pair of birth-zero diagrams, as stated in Corollary 4.15. Then,
$d_{\bullet}(Y,Y^{\prime})=\max_{i=1,\dots,N^{\prime}}\min\{|d_{i}-d^{\prime}_{i}|,\max\{d_{i}/2,d^{\prime}_{i}/2\}\}\leq\max_{i=1,\dots,N^{\prime}}|d_{i}-d^{\prime}_{i}|=\|\mathrm{DV}(Y)-\mathrm{DV}(Y^{\prime})\|_{\infty}$
is obvious. On the other hand,
$2\,d_{\bullet}(Y,Y^{\prime})=\max_{i=1,\dots,N^{\prime}}\min\{2\,|d_{i}-d^{\prime}_{i}|,\max\{d_{i},d^{\prime}_{i}\}\}\geq\max_{i=1,\dots,N^{\prime}}|d_{i}-d^{\prime}_{i}|=\|\mathrm{DV}(Y)-\mathrm{DV}(Y^{\prime})\|_{\infty}.$
Indeed, if $2\,d_{\bullet}(Y,Y^{\prime})$ is realized by $2\,|d_{i}-d_{i}^{\prime}|$ then the inequality is clear; if it is realized by, say, $d_{i}$, then $d_{i}\geq d_{i}^{\prime}$, so that $|d_{i}-d_{i}^{\prime}|=d_{i}-d_{i}^{\prime}\leq d_{i}$. ∎
Remark 4.19.
It is not hard to find examples which illustrate the the bounds in Theorem 4.12 are tight. Indeed, take $Y=\{(0,1)\}$, $Y^{\prime}=\{(0,\varepsilon)\}$ for some small $\varepsilon>0$. Then $d_{L}(Y,Y^{\prime})=1/2$, while the death vectorization distance is $1-\varepsilon$. Taking $\varepsilon\to 0$ gets arbitrarily close to realizing the upper bound. Likewise, it is easy to find an example showing that the lower bound is tight.
Remark 4.20.
It is shown in [35, 36] that if $Y$ and $Y^{\prime}$ are diagrams derived from degree-0 persistent homology of the Vietoris-Rips complexes of finite metric spaces $X$ and $X^{\prime}$, respectively, that
$d_{B}(Y,Y^{\prime})\leq\|\mathrm{DV}(Y)-\mathrm{DV}(Y^{\prime})\|_{\infty}\leq 2d_{GH}(X,X^{\prime}),$
where $d_{GH}$ is the Gromov-Hausdorff distance between metric spaces (see [9]). Indeed, this follows from the main stability result [36, Theorem 5] together with [35, Proposition 1.5] (from the extended arXiv version of the paper), which specializes to the birth-zero setting.
By the general Gromov-Hausdorff stability of persistent homology [14, Theorem 3.1], our Theorem 4.18 only yields the upper Gromov-Hausdorff bound
$d_{L}(Y,Y^{\prime})\leq 4d_{GH}(X,X^{\prime}),$
and is therefore suboptimal. On the other hand, our theorem is purely at the level of diagrams, and the example of Remark 4.19 shows that the bound cannot be improved at that level.
4.4 Coarse Non-Embeddability
Bubenik and Wagner [8] showed that persistence diagrams equipped with the bottleneck distance do not admit a coarse embedding into a Hilbert space. Motivated by their work, we strengthen their non-embeddability result by proving that even the space of birth-zero persistence diagrams–whether equipped with the bottleneck distance or the erosion distance–fails to admit a coarse embedding into a Hilbert space. This is established in Theorem 4.21, which not only offers an alternative proof of the result by Bubenik and Wagner  [8], but also expands upon it by extending the analysis to include the erosion and landscape distances.
Theorem 4.21 (Coarse Non-Embeddability in Hilbert Spaces).
The space $(\mathsf{PDgm}^{0},d_{\bullet})$ of birth-zero persistence diagrams, equipped with the distance $d_{\bullet}=d_{B}=d_{E}=d_{L}$, does not admit a coarse embedding into a Hilbert space. Consequently, the full spaces of persistence diagrams $(\mathsf{PDgm},d_{\bullet})$, with $d_{\bullet}=d_{B}$ or $d_{\bullet}=d_{E}=d_{L}$, also fail to admit such an embedding.
To prove Theorem 4.21, we show that the image $\ell^{\infty}_{\geq}$ of the death-vectorization embedding (see Remark 4.17) does not coarsely embed into a Hilbert space. This follows from Proposition 4.23 which implies that $\ell^{\infty}_{\geq}$ contains an isometric copy of every finite metric space. By a result of Mitra and Virk [37, Corollary 4.7.], the presence of such isometric copies constitutes an obstruction to coarse embeddability into a Hilbert space.
Remark 4.22.
The following proposition is similar to [45, Lemma 9], where Wagner considers the space of persistence diagrams equipped with the $p$-Wasserstein distance for $p>2$. In contrast, we work with the space $\ell_{\geq}^{\infty}$, and the constant $c$ in our proof is different.
Proposition 4.23.
Let $d\in\mathbb{N}$. Every finite subset of $(\mathbb{R}^{d},\|\cdot\|_{\infty})$ admits an isometric embedding into $(\ell_{\geq}^{\infty},\|\cdot\|_{\infty})$.
Proof.
Let $A=\{a^{1},a^{2},\dots,a^{n}\}$ be a finite subset of $(\mathbb{R}^{d},\|\cdot\|_{\infty})$. Choose a constant $c>0$ such that
$c>\max\{\|a^{i}\|_{\infty},|a^{i}_{k}-a^{j}_{k^{\prime}}|:1\leq i,j\leq n,1\leq k,k^{\prime}\leq d\}$
Define the map
$\varphi:A\to\ell^{\infty}_{\geq},\qquad a^{i}\mapsto(2c(d+2-k)+a^{i}_{k})_{k=1}^{d}$
Here, we abuse notation and write $(2c(d+2-k)+a^{i}_{k})_{k=1}^{d}$ to denote the sequence $(2c(d+2-k)+a^{i}_{k})_{k=1}^{d}$ extended by appending infinitely many zeros. We first claim that $\varphi(a^{i})\in\ell^{\infty}_{\geq}$ for each $a^{i}$. Observe that $a^{i}_{k}+2c\geq a^{i}_{k+1}$ for each $k$ because
$a^{i}_{k}+2c=a^{i}_{k}+c+c>0+c=c>a^{i}_{k+1}$
and it follows that $\varphi(a^{i})\in\ell^{\infty}_{\geq}$ since
$\varphi(a^{i})\in\ell^{\infty}_{\geq}\iff 2c(d+2-k)+a^{i}_{k}\geq 2c(d+2-(k+1))+a^{i}_{k+1}\iff a^{i}_{k}+2c\geq a^{i}_{k+1}$
Now we show that $\varphi$ is an isometry. Let $a^{i}$ and $a^{j}$ be arbitrary in $A$. Then
$\|\varphi(a^{i})-\varphi(a^{j})\|_{\infty}$ $=\max_{1\leq k\leq d}\left|2c(d+2-k)+a^{i}_{k}-2c(d+2-k)-a^{j}_{k}\right|$
$=\max_{1\leq k\leq d}\left|a^{i}_{k}-a^{j}_{k}\right|=\|a^{i}-a^{j}\|_{\infty},$
which concludes the proof. ∎
Corollary 4.24.
The space $(\ell^{\infty}_{\geq},\|\cdot\|_{\infty})$ does not admit a coarse embedding into a Hilbert space.
Proof.
We begin by noting that every finite metric space $(X=\{x_{1},x_{2},\dots,x_{d}\},d_{X})$ isometrically embeds into a finite subset of $(\mathbb{R}^{d},\|\cdot\|_{\infty})$ via the map $x\mapsto(d_{X}(x,x_{i}))_{i=1}^{d}$—this is known as the Kuratowski embedding [32] (see also [28, Remark 1.7]). Thus, Proposition 4.23 implies that $(\ell^{\infty}_{\geq},\|\cdot\|_{\infty})$ contains an isometric copy of every finite metric space. By [37, Corollary 4.7], this implies that $(\ell^{\infty}_{\geq},\|\cdot\|_{\infty})$ does not coarsely embed into any Hilbert space. ∎
Proof of Theorem 4.21.
By Corollary 4.24, the space $(\ell^{\infty}_{\geq},\|\cdot\|_{\infty})$ does not admit a coarse embedding into a Hilbert space. As the spaces $(\mathsf{PDgm}^{0},d_{B})$ and $(\ell^{\infty}_{\geq},\|\cdot\|_{\infty})$ are bi-Lipschitz equivalent via the death-vectorization map (Corollary 4.18), $(\mathsf{PDgm}^{0},d_{B})$ also fails to admit such an embedding. Since the bottleneck and erosion distances are equivalent on $\mathsf{PDgm}^{0}$ (Theorem 4.12), the same conclusion holds for $(\mathsf{PDgm}^{0},d_{E})$. Consequently, the full spaces $(\mathsf{PDgm},d_{B})$ and $(\mathsf{PDgm}^{0},d_{E})$ also do not admit coarse embeddings into a Hilbert space. ∎
Acknowledgements
We would like to thank Parker Edwards, Woojin Kim and Facundo Mémoli for useful conversations and feedback which helped us to improve the paper. This research was partially supported by NSF grant DMS–2324962.